### v2

# Bounds on the realizations of zero-nonzero patterns and sign conditions of polynomials restricted to varieties and applications

Abstract.
We prove upper bounds which are independent of the dimension of the ambient space, on the number of realizable zero-nonzero patterns as well as sign conditions (when the field of coefficients is ordered) of a finite set of polynomials $\mathcal{P}$ restricted to some algebraic subset $V$ of an affine or projective space. Our bounds depend on the number and the degrees of the polynomials in $\mathcal{P}$, as well as the degree of $V$ and the dimension of $V$, but are independent of the dimension of the space in which $V$ is embedded. This last feature of our bounds is useful in situations where the ambient dimension is much larger than $\dim V$.
We give several applications of our results. We generalize existing results on bounding the speeds of algebraically defined classes of graphs, bounding the ${\varepsilon}$-entropy of real varieties, as well as lower bounds in terms of the number of connected components for testing membership in semi-algebraic sets using algebraic computation trees. Motivated by quantum complexity theory we introduce a notion of relative rank (additive as well as multiplicative) in finite dimensional vector spaces and algebras relative to a fixed algebraic subset in the vector space or algebra – which generalizes the classical definition of ranks of tensors. We prove a very general lower bound on the maximum relative rank of finite subsets relative to algebraic subsets of bounded degree and dimension which is independent of the dimension of the vector space or algebra. We show how our lower bound implies a quantum analog of a classical lower bound result of Shannon for Boolean circuits – that almost all Boolean functions require (classical) circuits of size at least $\Omega(2^{n}/n)$ even in the presence of a quantum oracle given by an algebraic subset (of the group of unitary transformations) of fixed degree and dimension.
Key words and phrases:
Sign conditions, zero-nonzero patterns, Betti numbers, semi-algebraic sets, ranks, quantum circuits
2000 Mathematics Subject Classification:
Primary 68Q12; Secondary 14P10, 81P68
1. Introduction
Let $k$ be a field. For $x\in k$, we will denote
$\mathrm{sgn}(x)=\begin{cases}0\text{ if }x=0,\\ 1\text{ else},\end{cases}$
and for any set of polynomials $\mathcal{P}\subset k[X_{1},\ldots,X_{N}]$ we will call an element of $\{0,1\}^{\mathcal{P}}$ to be a zero-nonzero pattern on $\mathcal{P}$.
If $k$ is an ordered field we will denote for $x\in k$,
$\mathrm{sign}(x)=\begin{cases}0\text{ if }x=0,\\ 1\text{ if }x>0,\\ -1\text{ if }x<0,\end{cases}$
and for any set of polynomials $\mathcal{P}\subset k[X_{1},\ldots,X_{N}]$ we will call an element of $\{0,1,-1\}^{\mathcal{P}}$ to be a sign condition on $\mathcal{P}$.
1.1. Background and prior work
Bounds on the number (or other topological invariants of the realizations like the number of connected components etc.) of realizable zero-nonzero patterns or sign conditions of finite sets of polynomials have had many important applications in mathematics and theoretical computer science. To pick a few notable ones: they play a key role in bounding the number of point configurations as well as combinatorial types of polytopes in $\mathbb{R}^{d}$ [20], in bounding the number of geometric permutations induced by transversals to convex subsets of $\mathbb{R}^{d}$ [21], bounding the number (also called the speed) of algebraically defined classes of graphs [1, 38], in recent applications of the polynomial method to incidence combinatorics [11, 46, 47, 44, 41, 23], in proving lower bounds for testing membership in semi-algebraic sets using algebraic computation trees [12, 50, 19], in proving lower bounds on projective dimensions of graphs and span programs [36], bounding the ${\varepsilon}$-entropy of and volumes of tubes around algebraic varieties [9, 51], and a host of other applications (including some new ones introduced in this paper). The following two results maybe considered as starting points.
Theorem A.
[36, 25] Let $k$ be a field and $\mathcal{P}\subset k[X_{1},\ldots,X_{N}]_{\leq d}$ be a finite set of polynomials of degree at most $d$. Then, the number of realizable zero-nonzero patterns on $\mathcal{P}$ (i.e. elements of $\{0,1\}^{\mathcal{P}}$) is bounded by $(O(sd))^{N}$, where $s=\mathrm{card}(\mathcal{P})$. (We say a zero-nonzero pattern $\pi\in\{0,1\}^{\mathcal{P}}$ is realizable if there exists $x\in k^{N}$, such that $\mathrm{sgn}(P(x))=\pi(P)$ for all $P\in\mathcal{P}$.)
Theorem B.
[48] Let $k$ be an ordered field and $\mathcal{P}\subset k[X_{1},\ldots,X_{N}]_{\leq d}$ be a finite set of polynomials of degree at most $d$. Then, the number realizable sign conditions on $\mathcal{P}$ is bounded by $(O(sd))^{N}$, where $s=\mathrm{card}(\mathcal{P})$. (We say a sign condition $\sigma\in\{0,1,-1\}^{\mathcal{P}}$ is realizable if there exists $x\in k^{N}$, such that $\mathrm{sign}(P(x))=\sigma(P)$ for all $P\in\mathcal{P}$.)
The above theorems have been generalized in many different directions. For example, if $k=\mathrm{R}$ is a real closed field (for example the field $\mathbb{R}$ of real numbers), it is possible to obtain bounds on the higher Betti numbers of the realizations of sign conditions (see Theorem C below). Motivated by applications in discrete geometry (bounding the number of geometric permutations, as well as incidences), there was a need for more precise bounds on the number of realizable sign conditions: for example, upper bounds on the number of realizable sign conditions or even on the number of different connected components of the realizations restricted to an algebraic subset of the ambient space which may be of much smaller dimension (see for example [21] for one such application where the dimension of the algebraic subset plays a key role).
Before proceeding further we recall some definitions from algebraic geometry and introduce some useful notations.
1.2. Basic definitions and notation
Let $k$ denote an algebraically closed field. An algebraic subset $V$ of $k^{N}$ 111 In the scheme-theoretic language, $k^{N}$ is identified with the set of closed $k$-points of the $n$-dimensional affine scheme $\mathbb{A}_{k}^{N}$ [24]. It comes equipped with the Zariski topology. However, we will avoid schematic language in this paper and use $k^{n}$ and $\mathbb{A}_{k}^{N}$ interchangeably. Similar, remark applies to the projective scheme $\mathbb{P}_{k}^{N}$ as well. is the set of common zeros of a set of polynomials in $k[X_{1},\ldots,X_{N}]$.
An algebraic subset admits a decomposition (in the Zariski topology) into a finite union of irreducible ones which are called varieties. The dimension of a variety $V$, denoted $\dim V$, is the Krull dimension of its coordinate ring $k[V]=k[X_{1},\ldots,X_{n}]/I(V)$, where $I(V)$ is the defining ideal of $V$ i.e. $I(V)=\{f\in k[X_{1},\ldots,X_{N}]\;\mid\;f(x)=0,x\in V\}$. The degree of $V$, denoted $\deg V$, is equal to the number of points in the intersection $V\cap L$ for a generically chosen affine subspace $L$ of dimension $N-\dim V$ in $k^{N}$.
Finally, we define the degree, $\deg(V)$, of an algebraic set $V$ to be the sum of the degrees of its irreducible components, and the dimension, $\dim(V)$, of $V$ to be the maximum of the dimensions of its irreducible components.
A subset of $S\subset k^{N}$ is called a constructible subset if it is defined by a Boolean formula with atoms of the form $P=0$ for $P\in k[X_{1},\ldots,X_{N}]$. A subset of $S\subset\mathbb{P}_{k}^{N}$ is called a constructible subset if it is defined by a Boolean formula with atoms of the form $P=0$ for $P\in k[X_{0},X_{1},\ldots,X_{N}]$, with $P$ homogeneous.
If $\mathrm{R}$ is a real closed field (for example, $\mathbb{R}$), a subset of $S\subset\mathrm{R}^{N}$ is called a semi-algebraic subset if it is defined by a Boolean formula with atoms of the form $P=0,P>0$ for $P\in\mathrm{R}[X_{1},\ldots,X_{N}]$. We call a subset of $\mathbb{A}_{\mathrm{R}}^{N}$ or $\mathbb{P}_{\mathrm{R}}^{N}$ to be $\mathcal{P}$-semi-algebraic, for a finite set of polynomials $\mathcal{P}$, if every $P$ in the Boolean formula defining the set belongs to $\mathcal{P}$.
If $\mathrm{C}=\mathrm{R}[i]$ the algebraic closure of $\mathrm{R}$, then a constructible subset of $\mathrm{C}^{N}$ is also semi-algebraic subset of $\mathrm{R}^{2N}=\mathrm{C}^{N}$.
Notation 1.1.
In case, $k=\mathrm{C}=\mathrm{R}[i]$, where $\mathrm{R}$ is a real closed field, for any locally closed semi-algebraic set $S$ we denote by
$b_{i}(S)=\dim_{\mathbb{Z}/2\mathbb{Z}}\mbox{\rm H}_{i}(S,\mathbb{Z}/2\mathbb{Z}),$
where $\mbox{\rm H}_{i}(S,\mathbb{Z}/2\mathbb{Z})$ denotes the $i$-th homology group of $S$ with coefficients in $\mathbb{Z}/2\mathbb{Z}$ (see for example, [7, Chapter 6, Section 3]) and call $b_{i}(S)$ the $i$-th Betti number of $S$ (with coefficients in $\mathbb{Z}/2\mathbb{Z}$). Note that $b_{0}(S)$ is just the number of semi-algebraically connected components of $S$.  [^2] We say “semi-algebraically connected” rather than connected since $\mathrm{R}$ is assumed to be an arbitrary real closed field, possibly non-archimedean. “Semi-algebraically connected” means “semi-algebraically path connected” (see for example [7, Chapter 3, Section 2]), i.e. every two points in the set can be joined by a semi-algebraic path whose image is contained in the set. If $\mathrm{R}=\mathbb{R}$, then being semi-algebraically connected and being connected are equivalent properties of semi-algebraic sets.
We can now state several generalizations of the two basic theorems mentioned before. The following theorem extends Theorem B in several directions.
Theorem C.
[6] Let $\mathrm{R}$ be a real closed field and $V\subset\mathrm{R}^{N}$ be an algebraic subset defined by polynomials of degrees bounded by $d$ with $\dim_{\mathrm{R}}V=p$, and $\mathcal{P}\subset\mathrm{R}[X_{1},\ldots,X_{N}]_{\leq d}$. Then, for each $i,0\leq i\leq p$, the sum of the $i$-th Betti numbers of the realizations in $V$ of all sign conditions on $\mathcal{P}$ is bounded by
$s^{p-i}\cdot O(d)^{N},$
where $s=\mathrm{card}(\mathcal{P})$.
However, the above result is still insufficient for certain modern applications involving applications of the polynomial partitioning technique, where one needs more refined results where the dependence of the bound on the degree of variety $V$ and degrees of the polynomials in $\mathcal{P}$ are separated out.
Suppose that we have a sequence of algebraic sets $V_{0},\ldots,V_{\ell}\subset\mathrm{R}^{N}$ cut out by a sequence of real polynomials of degrees $d_{1},d_{2},\ldots,d_{\ell}$, with real dimensions $p_{1}\geq p_{2}\geq\cdots\geq p_{\ell}$, and the degrees of the polynomials in $\mathcal{P}$ are bounded by $d\geq d_{\ell}$.
The following result in this direction is proved in [4], and has been applied in the context of incidence geometry [11, 46, 47].
Theorem D.
[4, Theorem 4] Suppose that
$2\leq d_{1}\leq d_{2}\leq\frac{1}{N+1}d_{3}\leq\frac{1}{(N+1)^{2}}d_{4}\leq\cdots\leq\frac{1}{(N+1)^{\ell-2}}d_{\ell},$
and $d_{\ell}\leq\frac{1}{N+1}d$. Then, the number of semi-algebraically connected components of the realizations of all sign conditions on $\mathcal{P}$ restricted to $V=V_{\ell}$ is bounded by
$O(N)^{2N}(sd)^{p_{\ell}}\left(\prod_{1\leq j\leq\ell}d_{j}^{p_{j-1}-p_{j}}\right).$
The following related result which has applications in incidence geometry via the polynomial partitioning technique (see for example [23, 41] was obtained by by M. Walsh [47]. We denote by ${\rm Z}(P,V)$ the set of zeros of a polynomial restricted to a set $V$.
Theorem E.
[47, Theorem 2.12] Let $\mathrm{R}$ be a real closed field, $\mathrm{C}=\mathrm{R}[i]$ the algebraic closure of $\mathrm{R}$, and $V\subset\mathrm{C}^{N}$ be an algebraic variety with $\dim V=p$, and $P\in\mathrm{R}[X_{1},\ldots,X_{n}]$. Then the number of semi-algebraically connected components $C$ of $\mathrm{R}^{N}-{\rm Z}(P,\mathrm{R}^{N}))$ such that $C\cap V(\mathrm{R})\neq\emptyset$ is bounded by $C(N)\cdot\deg(V)\cdot\deg(P)^{p}$ where the constant $C(N)$ depends on $N$ (in an unspecified way), and $V(\mathrm{R})$ denotes the $\mathrm{R}$-points of $V$.
Remark 1.
Note that the quantity being bounded in Theorem E is a priori smaller than the number of semi-algebraically connected components of $V(\mathrm{R})-{\rm Z}(P,\mathrm{R}^{N})$. In particular, one cannot derive any useful bound on the number of semi-algebraically connected components of $V(\mathrm{R})$ itself from Theorem E (by letting $P$ to be a non-zero constant polynomial).
Remark 2.
Notice also that while Theorem C has better dependence in the combinatorial part (the part depending on $s$) on the dimension of the variety $V$, the algebraic part (the part depending on the degree) is oblivious of the dimension of the variety and depends exponentially on $N$ the dimension of the ambient space. The bounds in Theorems D and E are better but they still depend on $N$ (through the quantity $C=C(N)$, presumably exponential, in the case of Theorem E).
In many applications, the dimension of $V$ may be much smaller than $N$. For instance, consider the Plücker embedding of the Grassmannian variety $\mathrm{Gr}(m,M)$ of $m$-dimensional subspaces in $k^{M}$ in the $\binom{M}{m}-1$ dimensional projective space. It is well known that $\dim\mathrm{Gr}(m,M)=m(M-m)$. Thus, if $m=\lceil M/2\rceil$, while the dimension of the variety $V=\mathrm{Gr}(m,M)$ is polynomial in $M$, the dimension $N$ of the ambient projective space is exponentially large in $M$. In such a situation, the bounds in Theorems D and E which have exponential dependence on $N$, would produce a bound doubly exponential in $M$.
Laszlo and Viterbo [31] proved the following:
Theorem F.
[31, Theorem 1.2] The sum of all $\mathbb{Z}/2\mathbb{Z}$-Betti numbers of $V(\mathrm{R})$, where $V$ is a non-singular real variety of degree $D$ and dimension $p$, is bounded by
$2^{p^{2}+2}D^{p+1}.$
Remark 3.
The proof of Theorem F relies on sophisticated calculation of characteristic classes of the underlying smooth manifold of $V$ and hence cannot be applied to singular varieties.
Remark 4.
The bound in Theorem F in particular gives a bound on the number of semi-algebraically connected components of $V$ which just depends on the degree of $V$ and independent of the dimension of ambient space and is closest in spirit to the results in this paper. This bound was further improved to $D^{p+1}$ in [27]. However, these bounds only apply to non-singular real projective varieties – while in the results about sign conditions proved in this paper (see Theorem 2 and Remark 7 below) the hypothesis on $V$ is much weaker – namely that $V$ only needs to be a union of affine varieties $V_{i}$, such that each $V_{i}$ has the same real dimension at all points of $V_{i}(\mathrm{R})$. Many naturally occurring real varieties, such as Schubert subvarieties of flag varieties, rank subvarieties in spaces of matrices, as well as other determinantal varieties (see [33] for definitions), are often singular but satisfies the real equi-dimensionality condition mentioned above.
1.3. Summary of main contributions
In this paper we consider varieties and polynomials over real and complex numbers (in fact more generally over arbitrary real closed and algebraically closed fields). We prove bounds on the number of realizable zero-nonzero patterns and sign conditions (and also on the number of connected components of their realizations which is a priori bigger number) which are independent of $N$ and depend only on the number and degrees of the polynomials, and the degree (of the embedding of $V$ in $\mathbb{A}_{k}^{N}$ or in $\mathbb{P}_{k}^{N}$) and the dimension of $V$. These appear as Theorems 1, 2, 4, 5 and 6 in Section 2 below.
The independence of these bounds from the dimension $N$ of the ambient space is a new feature of these bounds. This feature is potentially useful in any application where one is interested in the number of zero-nonzero patterns (resp. sign conditions) restricted to algebraic subsets of dimension much smaller than that of the ambient dimension. We discuss examples of such applications below in Section 1.3.4.
1.3.1. Methods
Our techniques for dealing with zero-nonzero patterns and sign conditions are different. To prove the upper bound on the number of zero-nonzero patterns of a family of polynomials restricted to an algebraic subset $V$ of $\mathbb{A}_{k}^{N}$, we closely follow the proof given in [25, 8], with the modification that all algebraic subsets that are considered in the proof are contained in $V$ and we can bound their degrees in terms of $\deg V$ and the degrees of the given polynomials. We prove an upper bound on the sum of the degrees of the Zariski closure of the realizations of all zero-nonzero patterns with non-empty realizations on $V$, which in turn gives an upper bound on the number of realized zero-nonzero patterns. An important role in the argument is played by the generalized Bezout inequality (Theorem I).
For the problem of bounding realizable sign conditions over ordered fields we assume without loss of generality that the field $k=\mathrm{R}$ is a real closed field. The main technique that is used for proving bounds on the number of realizable sign conditions (or more generally bounding the Betti numbers of these realizations) is via a technique of infinitesimal perturbations (see for example, [6, 7]). The idea is to perturb the given set of polynomials using infinitesimal elements in some real closed extension of $\mathrm{R}$, so that:
  1. (a)
    The polynomials in the new set have the same degrees as ones in the given set, but in addition have better algebraic properties – like being in general position. These algebraic properties help control the combinatorial part of the bound.
  2. (b)
    The realizations of the various realizable sign conditions are related to those of the given set in some way, so that bounding the number (or more generally the Betti numbers) of realizations of the sign conditions of the new set of polynomials, suffices to bound the corresponding quantity for the original set.
  3. (c)
    Finally, the problem of bounding the number of realizable sign conditions (or more generally the Betti numbers of the realizations) can be reduced to bounding the corresponding quantity for all the non-empty algebraic subsets that one can define using the perturbed set of polynomials.
Our technique for proving upper bounds on the number of realizable sign conditions follow a similar paradigm but meets two important obstacles.
  1. (a)
    Since we are not given the polynomials that define $V$, we are not able to deform $V$ infinitesimally maintaining its degree;
  2. (b)
    as a result while we are able to reduce using usual techniques to the case bounding the number of semi-algebraically connected components of certain algebraic subsets of $V(\mathrm{R})$, all existing bounds for bounding the number of semi-algebraically connected components of possibly singular real algebraic sets depend exponentially on the dimension of the ambient space.
We overcome these problems by first reducing the problem to the problem of bounding the number of semi-algebraically connected components of an appropriate number of real algebraic subsets $X(\mathrm{R})\subset V(\mathrm{R})$, of co-dimension one in $V$. then after taking the image under a generic projection $\pi$ to $\mathrm{R}^{p}$ (where $p=\dim V$), to the problem of bounding the number of semi-algebraically connected components of the semi-algebraic set consisting of the $\mathrm{R}$-points of $\pi(X)-\mathrm{Sing}(\pi(X))$. We can use the classical bound due to Oleĭnik and Petrovskiĭ [34] to bound the latter quantity. This bound depends exponentially on the dimension of the ambient space, but since this is $p$ in this situation we can tolerate the exponential dependence on $p$. The technique described above for obtaining bounds on the number of semi-algebraically connected components of a real variety only in terms of its degree is new and in our opinion the most important technical contribution of the paper.
1.3.2. Tightness
Our bound (on the sum of the degrees of the realizations) in the algebraically closed case is tight (see Remark 5). In the real closed case, our main result applies to a certain slightly restricted class of algebraic subsets of $\mathrm{R}^{N}$ that includes most real algebraic sets that occur in practice. The bound is tight in this case as well up to a factor of $O(1)^{\dim V}$ (see Remark 8).
1.3.3. Connection with model theory and NIP structures
Bounding the number of realizable zero-nonzero patterns of $n$ subsets of a fixed definable set $V$ all coming from fixed definable family plays an important role in model theory in relation to an important class of theories called NIP theories. These are complete theories whose models have the property that every definable family of subsets has finite Vapnik-Chervonenkis dimension (see for example [35, Chapter 12, Section 4] or [42] for the precise definition of NIP theories and structures). A typical result (valid for instance over o-minimal theories, theories of algebraically closed fields, algebraically closed valued fields, which are all NIP) is that the number of realized zero-nonzero patterns on a fixed definable set $V$ realized by $n$ definable sets drawn from a definable family $\mathcal{H}$ is bounded by
$C_{V,\mathcal{H}}\cdot n^{\dim V},$
where $C_{V,\mathcal{H}}>0$ is a constant depending only on $V$ and $\mathcal{H}$ (see for example [3, 2, 5, 10] for such bounds in different theories). The exponent $\dim V$ is then an upper bound on the Vapnik-Chervonenkis codensity (see any of the papers mentioned above for definition) of the family $\mathcal{H}$. In fact, if $V$ belongs to some definable family $\mathcal{V}$, the constant, $C_{V,\mathcal{H}}$ can be made to depend just on $\mathcal{V},\mathcal{H}$. If we formulate Theorem 1 in the above language, the corresponding families $\mathcal{H}$ and $\mathcal{V}$ are as follows:
$\mathcal{H}=\mathcal{H}_{d}:=\bigcup_{N>0}\mathcal{H}_{d,N},$
where $\mathcal{H}_{d,N}$ consists of hypersurfaces of degree $d$ in $k^{N}$, and
$\mathcal{V}=\mathcal{V}_{D,p}:=\bigcup_{N>0}{V}_{D,p,N}$
where ${V}_{D,p,N}$ is the family of algebraic subsets of $k^{N}$ of degree bounded by $D$ and dimension $p$. Theorem 1 implies that the number of zero-nonzero patterns realized by any $n$ members of $\mathcal{H}$ on any member $V$ of $\mathcal{V}$ is bounded by $C_{\mathcal{V},\mathcal{H}}\cdot n^{\dim V}$, with $C_{\mathcal{V},\mathcal{H}}=C_{D,d,p}(=(O(Dd))^{p})$. This is similar to the prior results in model theory bounding zero-nonzero patterns mentioned above [5, 3, 2, 10], but notice that the families $\mathcal{H}_{d},\mathcal{V}_{D,p}$ are now countable unions of definable families and are not definable themselves. So the above mentioned implication of Theorem 1 cannot be deduced directly from the model theoretic bounds on zero-nonzero patterns (in the theory of algebraically closed fields). A similar remark applies to Theorem 2 as well.
1.3.4. Applications
We give several applications of our results. Some of these results improve or generalize prior results leveraging our new bounds. These include upper bounds on the “speeds” (see Definitions 3.1 and 3.2) of algebraically defined families graphs (Theorems 8 and 9), the ${\varepsilon}$-entropy of real varieties (Theorem 10), and lower bounds on the depths of algebraic computations trees testing membership in a semi-algebraic set (Theorem 11). In each of these cases the new bounds depend only on the degree and dimension of the relevant algebraic subset and independent of (or in the ${\varepsilon}$-entropy case very weakly dependent on) the ambient dimension. We also introduce a notion of additive (resp. multiplicative) rank of elements of a vector space (resp. algebra) relative to some algebraic subset $\Delta$ (Definitions 3.4 and 3.5). This is motivated from quantum complexity theory, where $\Delta$ can be thought of as an “quantum oracle” (i.e. a set of unitary operators operating on the Hilbert space associated to a $n$-qubit quantum register each having unit cost). We extend the definition of ranks to subsets of the vector space (resp. algebra), and prove lower bounds on the rank of subsets which have some structure (Theorems 12, 13, 14 and 15). As a consequence, finite subsets which are “grid”-like (i.e. has some algebraic structure) must contain elements of relatively large rank. As an application of the rank lower bound theorems we show how to recover Shannon’s lower bound theorem for classical circuits in the case of quantum circuits equipped with an algebraic oracle $\Delta$ (Theorems 16 and 17). In all these applications, the dimension and the degree of $\Delta$ is small compared to the dimension of the ambient vector space or algebra. So it is crucial that the bounds on the number of zero-nonzero patterns or sign conditions that are at the heart of these results be independent of the ambient dimension (for example, this dimension is exponential in the number of qubits in the application to quantum complexity).
The rest of the paper is organized as follows. We state our mathematical results in the following Section 2. In Section 3 we describe the applications of theorems stated in Section 2. These include:
  1. (a)
    bounding speeds of algebraically defined families of graphs (Section 3.1);
  2. (b)
    bounding the ${\varepsilon}$-entropy of real varieties (Section 3.2);
  3. (c)
    an improvement of known lower bound results for algebraic computation trees (Section 3.3);
  4. (d)
    lower bounds on the relative ranks, additive as well as multiplicative in vector spaces and algebras respectively, and over algebraically closed as well as real closed fields (Section 3.4); and finally,
  5. (e)
    how the lower bound on ranks of subsets in an algebra imply lower bounds on quantum circuit complexity with oracle (Section 3.5).
Section 4 is devoted to the proofs of all the results mentioned above.
2. Upper bounds on zero-nonzero patterns and sign conditions
2.1. Bound on the number of realizable zero-nonzero patterns
For any finite set $\mathcal{P}$ of polynomials in $k[X_{1},\ldots,X_{N}]$, $V$ an algebraic subset of $k^{N}$, and $\pi\in\{0,1\}^{\mathcal{P}}$, we denote by
${\mathcal{R}}(\pi,V)=\{x\in V\mid\mathrm{sgn}(P(x))=\pi(P),P\in\mathcal{P}\}.$
We will denote
$\Pi(\mathcal{P},V)=\{\pi\in\{0,1\}^{\mathcal{P}}\;\mid\;{\mathcal{R}}(\pi,V)\neq\emptyset\}$
and call $\Pi(\mathcal{P},V)$ the set of realizable zero-nonzero patterns of $\mathcal{P}$ on $V$.
We are now in a position to state our first result.
Theorem 1.
Let $V\subset k^{N}$ (resp. $V\subset\mathbb{P}_{k}^{n}$) be an algebraic subset with $\deg V=D$, $\dim V=p$, and $\mathcal{P}\subset k[X_{1},\ldots,X_{N}]_{\leq d}$ (resp. $\mathcal{P}\subset k[X_{0},X_{1},\ldots,X_{N}]_{\leq d}$) be a finite set of polynomials (resp. homogeneous polynomials), with $\mathrm{card}(\mathcal{P})=s$. Then,
$\mathrm{card}(\Pi(\mathcal{P},V))\leq\sum_{\pi\in\{0,1\}^{\mathcal{P}}}\deg\overline{({\mathcal{R}}(\pi,V))}\leq D\cdot\sum_{i=0}^{p}\binom{s}{i}\cdot d^{i}$
(where $\overline{({\mathcal{R}}(\pi,V))}$ denotes the Zariski-closure of ${\mathcal{R}}(\pi,V)$).
Remark 5 (Tightness).
The bound in (the second inequality of) Theorem 1 is tight. To see this let $N>2p$, and $V$ be a union of $D$ disjoint affine subspace of dimension $p$ in $k^{N}$. Let $\mathcal{P}$ be a set of $s$ polynomials in $k[X_{1},\ldots,X_{N}]$, where each $P\in\mathcal{P}$ is a product of $d$ generic polynomials of degree $1$. It is easy to see that with these choices,
$\sum_{\pi\in\{0,1\}^{\mathcal{P}}}\deg\overline{({\mathcal{R}}(\pi,V))}=D\cdot\sum_{i=0}^{p}\binom{s}{i}\cdot d^{i}.$
2.2. Bound on the number of realizable sign conditions
Now suppose $\mathrm{R}$ is a real closed field and let $\mathrm{C}=\mathrm{R}[i]$. For an algebraic subset $V\subset\mathrm{C}^{N}$, we denote by $V(\mathrm{R})=V\cap\mathrm{R}^{N}$ the set of its real points. We say that $V$ is a real algebraic set (or a real variety if $V$ is irreducible), if $V$ is closed under conjugation taking $z$ to $\bar{z}$. It is easy to see that a real algebraic set is the vanishing locus of a finite number of polynomials in $\mathrm{R}[X_{1},\ldots,X_{N}]$.
For $x\in V(\mathrm{R})$, we will denote by $\dim_{x}V(\mathrm{R})$ to be the local real dimension (see for example [7, page 191]) of $V(\mathrm{R})$ at $x$ (as a semi-algebraic set). If $V$ is a non-singular variety of dimension $p$, then $\dim_{x}V(\mathrm{R})=p$ for every $x\in V(\mathrm{R})$.
For any finite set $\mathcal{P}$ of polynomials in $\mathrm{R}[X_{1},\ldots,X_{N}]$, $S$ a semi-algebraic subset of $\mathrm{R}^{N}$, and $\sigma\in\{0,1,-1\}^{\mathcal{P}}$, we denote by
${\mathcal{R}}(\sigma,S)=\{x\in S\mid\mathrm{sign}(P(x))=\sigma(P),P\in\mathcal{P}\}.$
We will denote
$\Sigma(\mathcal{P},S)=\{\sigma\in\{0,1,-1\}^{\mathcal{P}}\;\mid\;{\mathcal{R}}(\sigma,S)\neq\emptyset\},$
and call $\Sigma(\mathcal{P},S)$ the set of realizable sign conditions of $\mathcal{P}$ on $S$.
Notation 2.1.
For $S$ a semi-algebraic set, we will denote by $\mathrm{Cc}(S)$ the set of semi-algebraically connected components of $S$.
Remark 6.
Note that $\mathrm{card}(\mathrm{Cc}_{0}(S))=b_{0}(S)$.
We can now state our main theorems on bounding the number of semi-algebraically connected components of the realizations of realizable sign conditions of a finite set of polynomials restricted to the real points of a given algebraic set $V$.
Theorem 2.
Let $V=V_{1}\cup\cdots\cup V_{m}\subset\mathrm{C}^{N}$ be a union of real varieties, with $\deg(V)=D$, $\dim V=p$, and such that for each $i,1\leq i\leq m$, and $x\in V_{i}(\mathrm{R})$, $\dim_{x}V_{i}(\mathrm{R})=\dim V_{i}$. Let $\mathcal{P}\subset\mathrm{R}[X_{1},\ldots,X_{N}]_{\leq d}$ be a finite set of polynomials, with $\mathrm{card}(\mathcal{P})=s$. Then,
$\mathrm{card}(\Sigma(\mathcal{P},V))\leq\sum_{\sigma\in\{0,1,-1\}^{\mathcal{P}}}b_{0}({\mathcal{R}}(\sigma,V(\mathrm{R})))$$\leq$ $O(D)^{p}\cdot\left(\left(\sum_{j=1}^{p}4^{j}\binom{s}{j}d^{p}\right)+D\right)$
$\leq$ $O(D)^{p}((sd)^{p}+D).$
Remark 7.
The hypothesis on $V$ is clearly satisfied if each $V_{i}$ is non-singular of dimension $p$ (see also Remark 4).
Remark 8 (Tightness).
Consider the following example taking $p=1$ and the dimension of the ambient space $N=2$.
Let $V$ be the algebraic subset of $\mathrm{C}^{2}$ defined by $Q=0$, where
$Q=\sum_{i=1}^{2}\prod_{j=0}^{D-1}(X_{i}-j)^{2}-{\varepsilon},$
where ${\varepsilon}>0$ and will be chosen later. Then, $p=\dim V=1$, and $\deg V=2D$. Moreover, for all small enough ${\varepsilon}>0,{\varepsilon}\in\mathrm{R}$, $V$ is non-singular and $V(\mathrm{R})$ consists of $D^{2}$ disjoint ovals, $V_{i,j},0\leq i,j,\leq D-1$, where $V_{i,j}$ is a oval with the point $(i,j)$ in its interior.
Let $\mathcal{P}=\{P_{1},\ldots,P_{s}\}$ be the set of polynomials, where
$P_{i}=\prod_{j=1}^{d}(X_{1}-\delta_{i,j}).$
Then for
$0<\delta_{1,1}\ll\delta_{1,2}\ll\cdots\ll\delta_{1,d}\ll\delta_{2,1}\ll\cdots\ll\delta_{s,d}\ll{\varepsilon}\ll 1,$
(where $\ll$ means sufficiently small), it is easy to verify that for each $i,1\leq i\leq s$, ${\rm Z}(P_{i},\mathrm{R}^{2})$ consists of $d$ vertical lines which meets each of the ovals $V_{0,0},\ldots,V_{0,D-1}$ in $2d$ distinct points.
Thus, ${\rm Z}(\mathcal{P},\mathrm{R}^{2})$ meets each of the ovals $V_{0,0},\ldots,V_{0,D-1}$ is $2sd$ distinct points, and the complement of these points in each oval is a set of $2sd$ disjoint arcs. Hence,
$\sum_{\sigma\in\{0,1,-1\}^{\mathcal{P}}}b_{0}({\mathcal{R}}(\sigma,V(\mathrm{R})))$ $=$ $(4sdD+D(D-1))$
$=$ $D(4sd+D-1)$
$=$ $\Omega(\deg V)((sd)+\deg V),$
matching the upper bound in Theorem 2 with $p=1$. Similar examples can be created for $p>1$ as well.
Theorem 2 may be interesting even in the case $\mathcal{P}=\emptyset$. In this case we just obtain a bound on the number of semi-algebraically connected components of $V$ having the top real dimension. In order to state this formally we introduce a new notation.
Notation 2.2.
For any semi-algebraic set $S$ and $p\geq 0$, we denote
$S_{p}=\{x\in S\mid\dim_{x}S\geq p\}.$
We have the following slight variant of Theorem 2 in the case of algebraic sets that will be useful later.
Theorem 3.
Let $V\subset\mathrm{C}^{N}$ be a real algebraic set, with $\deg(V)=D$, and $\dim V=p$. Then,
(2.1) $b_{0}(V(\mathrm{R})_{p})\leq 2^{2p+3}\cdot D^{p+1}=(O(D))^{p+1}.$
In particular, if every semi-algebraically connected component of $V(\mathrm{R})$ has (real) dimension equal to $p$, then
$b_{0}(V(\mathrm{R}))\leq 2^{2p+3}\cdot D^{p+1}=(O(D))^{p+1}.$
Remark 9.
As remarked earlier, many naturally occurring real varieties, such as Schubert subvarieties of flag varieties, rank subvarieties in spaces of matrices, as well as other determinantal varieties, are not non-singular but satisfies the hypothesis of Theorem 2. These varieties in general will not satisfy the hypothesis of Theorem F because of the presence of singularities.
However, there exist real varieties for which the hypothesis of Theorem 2 is not satisfied.
  1. (a)
    The singular curve $V\subset\mathrm{C}^{2}$ in the plane defined by $Y^{2}=X^{2}(X-1)$. In this case the $\dim V=1$, but $V(\mathrm{R})$ contains the isolated point $x=(0,0)$, and $\dim_{x}V(\mathrm{R})=0<1$. Thus, $V$ does not satisfy the hypothesis of Theorem 2.
  2. (b)
    Another interesting example is the well-known Whitney umbrella, $V\subset\mathrm{C}^{3}$, defined by the equation $X^{2}-ZY^{2}=0$. In this case $V(\mathrm{R})$ has only one connected component but every point $x\in V(\mathrm{R})$ on the “stick” of the umbrella (namely the points satisfying $Z<0$) has local dimension $\dim_{x}V(\mathrm{R})=1<2=\dim V$. The real variety $V$ does not satisfy the hypothesis of Theorem 2 (because of the lower dimensional stick of the umbrella).
We also prove a projective version of Theorem 2. It is no longer meaningful to talk about the sign of a (homogeneous) polynomial in a real projective space. However, one still has the notion of zero-nonzero patterns. We bound the number of semi-algebraically connected components of the realizations on $V(\mathrm{R})$ of all zero-nonzero patterns of a finite set of homogeneous polynomials, where $V$ is a non-singular real projective variety.
Theorem 4 (Projective case).
Let $V=V_{1}\cup\cdots\cup V_{m}\subset\mathbb{P}_{\mathrm{C}}^{N}$ be a union of real varieties, with $\deg(V)=D$, $\dim V=p$, and such that for each $i,1\leq i\leq m$, and $x\in V_{i}(\mathrm{R})$, $\dim_{x}V_{i}(\mathrm{R})=\dim V_{i}$, and let $\mathcal{P}\subset\mathrm{R}[X_{0},\ldots,X_{N}]_{\leq d}$ be a finite set of homogeneous polynomials, with $\mathrm{card}(\mathcal{P})=s$. Then,
$\mathrm{card}(\Pi(\mathcal{P},V(\mathrm{R})))\leq\sum_{\pi\in\{0,1\}^{\mathcal{P}}}b_{0}({\mathcal{R}}(\pi,V(\mathrm{R})))\leq O(D)^{p}((sd)^{p}+D).$
Remark 10.
Aside from the other applications that we discuss in the next section Theorem 4 can be used to simplify proofs of certain quantitative results in discrete geometry. For example in [21], the authors prove a bound on the number of geometric permutations induced by $k$-dimensional transversals to a finite family of convex sets in $\mathrm{R}^{d}$. A key role is played by a bound on the number of realizable sign conditions of a certain finite set of polynomials restricted to the Grassmannian $\mathrm{Gr}(k,d)$ (more accurately, the affine Grassmannian but this is unimportant). In order to get a tight bound on this quantity that does not depend on the dimension of the Plücker emdedding the authors utilize a fact from real algebraic geometry – that the Grassmannian varieties, $\mathrm{Gr}(m,N)$ can be embedded semi-algebraically (not scheme theoretically) as an algebraic subset of $\mathrm{R}^{N^{2}}$ [13, Theorem 3.4.4]. This last fact is somewhat remarkable as there is no such affine embedding of the complex Grassmannian (indeed the complex Grassmannian is compact as a topological space, while every complex affine algebraic set of positive dimension is necessarily non-compact). However, using Theorem 4 (noting that the Grassmannians are all non-singular projective varieties) and the known expressions for the degree and the dimension of $\mathrm{Gr}(m,N)$ (see (3.5)) one can obtain a similar bound as in [21, Theorem 2] without resorting to the clever non-standard embedding of the real Grassmannian, and the new bound is a slight improvement over the original result when $k$ is close to $d$.
The bound in Theorem 2 is worse than in Theorem 1 (in the dependence on $D$ and $d$). We can prove a bound with a better dependence on $D$ but with a more stringent assumption on $V$.
Theorem 5.
Let $\overline{V}\subset\mathbb{P}_{\mathrm{C}}^{N}$ be a non-singular complete intersection real variety with $\deg\overline{V}=D$ and $\dim\overline{V}=p$. Let $V=\overline{V}\cap\mathrm{C}^{N}$ denote the corresponding affine variety.
Let $\mathcal{P}\subset\mathrm{R}[X_{1},\ldots,X_{N}]_{\leq d}$ be a finite set of polynomials with $d\geq D$ and $\mathrm{card}(\mathcal{P})=s$.
Then, for each $i,0\leq i\leq p$,
$\sum_{\sigma\in\{0,1,-1\}^{\mathcal{P}}}b_{i}({\mathcal{R}}(\sigma,V(\mathrm{R})))\leq s^{p-i}\cdot D^{2}\cdot(O(d))^{p}.$
In particular,
$\sum_{\sigma\in\{0,1,-1\}^{\mathcal{P}}}b_{0}({\mathcal{R}}(\sigma,V(\mathrm{R})))\leq D^{2}\cdot(O(sd))^{p}.$
We are able to remove the non-singularity restriction in Theorem 5. However, the bound that we prove is on a priori smaller quantity which nevertheless arises in applications (see for example Theorem E which plays an important role in applications of polynomial partitioning technique in problems of incidence geometry).
Notation 2.3.
For $S,V\subset\mathrm{R}^{N}$ semi-algebraic subsets, we will denote
$\mathrm{Cc}(S,V)=\{C\in\mathrm{Cc}(S)\mid C\cap V\neq\emptyset\}.$
(Note that $\mathrm{Cc}(S)=\mathrm{Cc}(S,\mathrm{R}^{N})$.)
Theorem 6.
Let $V\subset\mathrm{C}^{N}$ be a real complete intersection variety (not necessarily non-singular) with $\deg(V)=D$ and $\dim(V)=p$. Let $\mathcal{P}\subset\mathrm{R}[X_{1}\ldots,X_{N}]_{\leq d}$ be a finite set of polynomials, with $d\geq D$ and $\mathrm{card}(\mathcal{P})=s$.
Then,
$\sum_{\sigma\in\{0,1,-1\}^{\mathcal{P}}}\mathrm{card}(\mathrm{Cc}({\mathcal{R}}(\sigma,\mathrm{R}^{N}),V(\mathrm{R})))\leq D^{4}\cdot(O(sd))^{2p}.$
2.3. Sum of Betti numbers of realizations of sign conditions
Using the result of Laszlo and Viterbo (Theorem F) we prove the following theorem which extends Theorem C to the case when the ambient variety is more general than $\mathrm{R}^{N}$.
Theorem 7.
Let $\overline{V}\subset\mathbb{P}_{\mathrm{C}}^{N}$ be a non-singular real variety with $\deg V=D$ and $\dim V=p$. Let $V=\overline{V}\cap\mathrm{C}^{N}$ denote the corresponding affine variety. Let $\mathcal{P}\subset\mathrm{R}[X_{1},\ldots,X_{N}]_{\leq d}$ be a finite set of polynomials. Then, for each $i,0\leq i\leq p$,
$\sum_{\sigma\in\{0,1,-1\}^{\mathcal{P}}}b_{i}({\mathcal{R}}(\sigma,V(\mathrm{R})))\leq O(1)^{p}\cdot s^{p-i}\cdot D^{p}\cdot(D+d^{p}).$
Remark 11.
Notice that the hypothesis on the variety $V$ is much stronger than in Theorem C or Theorem 2. However, the bound unlike Theorem C is independent of the ambient dimension.
3. Applications
3.1. Speeds of algebraic families of graphs
Let $k$ be an algebraically closed field and suppose that $F\subset k^{n}\times k^{n}$ be a fixed constructible relation (i.e. a constructible subset of $k^{n}\times k^{n}$). A labelled (directed) graph $(V,E\subset V\times V)$, with $V=\{1,\ldots,n\}$ is called an $F$-graph if there exists $(\mathbf{y}_{1},\ldots,\mathbf{y}_{n})\in(k^{N})^{n}$, such that
$(i,j)\in E\Leftrightarrow(\mathbf{y}_{i},\mathbf{y}_{j})\in F.$
We call the family of $F$-graphs for a fixed constructible relation $F$, a constructible family of graphs.
Now suppose that $\mathrm{R}$ is a real closed field and suppose that $F\subset\mathrm{R}^{n}\times\mathrm{R}^{n}$ be a fixed semi-algebraic relation (i.e. a semi-algebraic subset of $\mathrm{R}^{n}\times\mathrm{R}^{n}$). A labelled (directed) graph $(V,E\subset V\times V)$, with $V=\{1,\ldots,n\}$ is called an $F$-graph if there exists $(\mathbf{y}_{1},\ldots,\mathbf{y}_{n})\in(\mathrm{R}^{N})^{n}$, such that
$(i,j)\in E\Leftrightarrow(\mathbf{y}_{i},\mathbf{y}_{j})\in F.$
We call the family of $F$-graphs for a fixed semi-algebraic relation $F$, a semi-algebraic family of graphs.
Families of $F$-graphs, where $F$ is a semi-algebraic relation, have been studied in graph theory from several perspectives such as Ramsey theory, Zarankiewicz problem etc. (see for example, [1, 17]). We will be concerned with another very well-studied property, called “speed”, of such families.
The function $G_{F}:\mathbb{N}\rightarrow\mathbb{N}$, defined by
$G_{F}(n)=\mbox{ number of distinct $F$-graphs with vertex set $[1,n]$}$
is called the speed of the family of graphs defined by $F$.
The following theorem (slightly simplified) is a key result in the theory of semi-algebraic graphs.
Theorem G.
[38, Theorem 1.2] For every semi-algebraic relation $F\subset\mathrm{R}^{N}\times\mathrm{R}^{N}$,
$G_{F}(n)\leq n^{(1+o_{F}(1))Nn}$
(where the $o_{F}(1)$ term depends on $F$).  [^3] Theorem 1.2 in [38] is slightly more general in that it gives a bound on the number of edge-labellings taking values in some fixed finite set $\Lambda$. The Theorem stated here is the special case when $\Lambda=\{0,1\}$.
Now suppose that $k$ is an algebraically closed field. Let $F\subset k^{N}\times k^{N})_{n>0}$ (resp. $F\subset\mathbb{P}_{k}^{N}\times\mathbb{P}_{k}^{N}$) be a constructible subset, and $Z\subset k^{N}$ (resp. $Z\subset\mathbb{P}_{k}^{N}$ ) an algebraic subset. We will call a labelled (directed) graph $G=(V,E\subset V\times V)$, with $V=[1,n]:=\{1,\ldots,n\}$ is called a $(Z,F)$-graph if there exists $(\mathbf{y}_{1},\ldots,\mathbf{y}_{n})\in Z^{n}$, such that
$(i,j)\in E\Leftrightarrow(\mathbf{y}_{i},\mathbf{y}_{j})\in F,$
and denote the set of $(Z,F)$-graphs with vertex set $[1,n]$ by $\mathcal{G}_{Z,F}(n)$.
Definition 3.1.
Let $\mathbf{F}=(F_{n}\subset k^{N_{n}}\times k^{N_{n}})_{n>0}$ (resp. $\mathbf{F}=(F_{n}\subset\mathbb{P}_{k}^{N_{n}}\times\mathbb{P}_{k}^{N_{n}})_{n>0}$) a sequence of constructible subsets, and $\mathbf{Z}=(Z_{n}\subset k^{N_{n}})_{n>0}$ (resp. $\mathbf{Z}=(Z_{n}\subset\mathbb{P}_{k}^{N_{n}})_{n>0}$) a sequence of algebraic subsets.
We denote
$G_{\mathbf{Z},\mathbf{F}}(n)=\mathrm{card}(\mathcal{G}_{Z_{n},F_{n}}(n)).$
(Note that if $\mathbf{F}=(F_{n}=F\subset k^{N}\times k^{N})_{n>0}$, and $\mathbf{Z}=(Z_{n}=k^{N})_{n>0}$, then $G_{\mathbf{Z},\mathbf{F}}=G_{F}$).
We prove (with the same notation as above):
Theorem 8.
Let $c_{1},c_{2},c_{3},p>0$ be fixed constants, $\mathbf{Z},\mathbf{F}$ be as in Definition 3.1, and moreover suppose that each $F_{n}$ is a $\mathcal{P}_{n}$-constructible subset for some finite set of polynomials $\mathcal{P}_{n}$ in $2N_{n}$-variables, with $\mathrm{card}(\mathcal{P})\leq c_{1}$, $\deg(P)\leq c_{2}n$ for all $P\in\mathcal{P}_{n}$, $\deg(Z_{n})\leq c_{3}n$, and $\dim Z_{n}\leq p$. Then,
$G_{\mathbf{Z},\mathbf{F}}(n)\leq n^{(1+o(1))(p+1)n}.$
Example 3.1.
For example, we can take (as a randomly chosen example satisfying the hypothesis of Theorem 8) $k=\mathbb{C}$, $N_{n}=n$, $Z_{n}\subset\mathbb{P}_{\mathbb{C}}^{n}$ a $K3$-surface in $\mathbb{P}_{\mathrm{C}}^{n}$ with $\deg(Z_{n})=2n-2$ (there is a $19$-dimensional family of these [22, page 593]), and $F_{n}\subset\mathbb{P}_{\mathbb{C}}^{n}\times\mathbb{P}_{\mathbb{\mathrm{C}}}^{n}$ defined by $H_{n}=0$, where $H_{n}=\sum_{0\leq i,j\leq n}X_{i}^{n}Y_{j}^{n}$. In this example, $\dim Z_{n}=2$, $\deg Z_{n}=2n-2$, and $F_{n}$ is $\{H_{n}\}$-constructible with $\deg(H_{n})=2n$. So using the notation of Theorem 8, $c_{1}=1$, $c_{2}=c_{3}=2$, and $p=2$.
Now suppose that $\mathrm{R}$ is a real closed field.
Definition 3.2.
Let $\mathbf{F}=(F_{n}\subset k^{N_{n}}\times k^{N_{n}})_{n>0}$ (resp. $\mathbf{F}=(F_{n}\subset\mathbb{P}_{\mathrm{R}}^{N_{n}}\times\mathbb{P}_{\mathrm{R}}^{N_{n}})_{n>0}$) a sequence of semi-algebraic subsets, and $\mathbf{Z}=(Z_{n}\subset\mathrm{R}^{N_{n}})_{n>0}$ (resp. $\mathbf{Z}=(Z_{n}\subset\mathbb{P}_{\mathrm{R}}^{N_{n}})_{n>0}$) a sequence of algebraic subsets.
As before, we will call a labelled (directed) graph $(V,E\subset V\times V)$, with $V=\{1,\ldots,n\}$ is called a $(\mathbf{Z},\mathbf{F})$-graph if there exists $(\mathbf{y}_{1},\ldots,\mathbf{y}_{n})\in Z_{n}^{n}$, such that
$(i,j)\in E\Leftrightarrow(\mathbf{y}_{i},\mathbf{y}_{j})\in F_{n}.$
The function $G_{\mathbf{Z},\mathbf{F}}:\mathbb{N}\rightarrow\mathbb{N}$ is defined as before.
We prove (with the same notation as above):
Theorem 9.
Let $c_{1},c_{2},c_{3},p>0$ be fixed constants, $\mathbf{Z},\mathbf{F}$ be as in Definition 3.2, and moreover suppose that each $F_{n}$ is a $\mathcal{P}_{n}$-semi-algebraic subset for some finite set of polynomials $\mathcal{P}_{n}$ in $2N_{n}$-variables, with $\mathrm{card}(\mathcal{P})\leq c_{1}$, $\deg(P)\leq c_{2}n$ for all $P\in\mathcal{P}_{n}$, and suppose that each $Z_{n}$ is a finite union of real complete intersection varieties, with $\deg(Z_{n})\leq c_{3}n$, and $\dim Z_{n}\leq p$. Then,
$G_{\mathbf{Z},\mathbf{F}}(n)\leq n^{O(pn)}.$
Remark 12.
Note that Theorems 8 and 9 generalizes Theorem G in several different ways. Unlike in Theorem G, the embedding dimension of the underlying constructible/semi-algebraic family of graphs, as well as the degrees of the polynomials and the algebraic subsets from which the points representing the vertices are chosen are all allowed to grow with $n$ (and the embedding dimension $N_{n}$ is allowed to depend arbitrarily on $n$ and be unbounded as a function of $n$). Because of this flexibility, Theorems 8 and 9 cannot be deduced (unlike Theorem G) from prior results on bounding the number of zero-nonzero patterns (Theorem A) or sign conditions (Theorems B and C).
A particular case might be of interest in the theory of geometrically defined graphs.
Example 3.2 (Schubert graphs).
Let $Z=\mathrm{Gr}(m,N)\subset\mathbb{P}_{k}^{\binom{N}{m}-1}$ denote the Plücker embedding of the Grassmannian of $m$-dimensional subspaces of $k^{N}$, and let $F$ be any constructible relation in $\mathbb{P}_{k}^{\binom{N}{m}-1}\times\mathbb{P}_{k}^{\binom{N}{m}-1}$. For example, the restriction of $F$ to $Z\times Z$ could be a “Schubert-type” relation – namely $(p,q)\in F$ if and only $\dim p\cap q\geq i$ for some fixed $i\geq 0$.
Then,
(3.1) $G_{Z,F}(n)=n^{(1+o(1))(m(N-m)+1)n}.$
Remark 13.
Notice that in Example 3.2 if one applies directly Theorem G then one would obtain
(3.2) $G_{Z,F}(n)\leq n^{(1+o(1))(\binom{N}{m}-1)n}$
which is considerably worse in the exponent for $m\sim N/2$. However, representing $m$-dimensional subspaces in $k^{N}$, using $N\times m$ matrices of full rank, it is possible to apply the same techniques as in the proof of Theorem G to get
(3.3) $G_{Z,F}(n)\leq n^{(1+o(1))mNn}$
which is better than inequality (3.2) but still worse than (3.1).
3.2. Bound on the ${\varepsilon}$-entropy of real varieties
In this section we assume that $\mathrm{R}=\mathbb{R}$.
A quantitative result from real (or more accurately metric) algebraic geometry that is now of great interest in applications (in problems originating in machine learning) is a tight upper bound on the volume of an ${\varepsilon}$-tube around a real algebraic variety $V$ intersected with a ball of a given radius. A recent result in this direction is [9, Theorem 1.1] which gives a tight upper bound on this volume depending on the degree $d$ of the polynomials defining $V$, the codimension of $V$.
More recently, using very different techniques (originally due to Vitushkin [45]), Zhang and Kileel [51] proved another bound on the volume of such tubes, which matches the bound in [9] in the regime large $N$ and fixed $d$. They applied their results to prove new bounds in approximation properties of low rank tensors, on the required dimension for randomized sketching algorithms, and generalization error bounds for deep neural networks with rational or ReLU activation functions. Their technique depends crucially on a bound on the number of connected components of real algebraic varieties and any improvement on this bound will automatically improve the bound in their paper. In particular, inequality (2.1) is applicable in their context.
We follow the same notation as in [51]. Let $V(\mathrm{R})\subset\mathbb{R}^{N}$ be a real variety, and for ${\varepsilon}>0$ denote by $\mathcal{N}(V(\mathrm{R}),{\varepsilon})$ the minimum number of ${\varepsilon}$-balls required to cover $V(\mathrm{R})\cap B_{N}(\mathbf{0},1)$. The quantity $\log\mathcal{N}(V(\mathrm{R}),{\varepsilon})$ is referred to as the ${\varepsilon}$-entropy of $V(\mathrm{R})\cap B_{N}(\mathbf{0},1)$ and an upper bound on $\mathcal{N}(V(\mathrm{R}),{\varepsilon})$ implies an upper bound on the volume of the ${\varepsilon}$-tube around $V(\mathrm{R})\cap B_{N}(\mathbf{0},1)$.
The following result appears in [51, page 4, Table 1, entry on real varieties].
Theorem H.
Let $V(\mathrm{R})\subset B_{N}(\mathbf{0},1)$ be a real variety defined by polynomials of degrees at most $d$ and $\dim V=p$. Then,
(3.4) $\log\mathcal{N}(V(\mathrm{R}),{\varepsilon})\leq p\cdot\log(1/{\varepsilon})+O(N\cdot\log d+p\cdot\log N).$
We prove the following theorem.
Theorem 10.
Let $V\subset\mathrm{C}^{N}$ be a real variety with $\deg(V)=D,\dim V=p$, such that $V(\mathrm{R})\subset B_{N}(\mathbf{0},1)$. Then,
$\log\mathcal{N}(V(\mathbb{R})_{p},{\varepsilon})\leq p\cdot\log(1/{\varepsilon})+O(p\cdot(\log\deg(V)+\log N))$
(see Notation 2.2).
In particular, if every point of $V(\mathrm{R})$ has local real dimension equal to $p$, then
$\log\mathcal{N}(V(\mathrm{R}),{\varepsilon})\leq p\cdot\log(1/{\varepsilon})+O(p\cdot(\log\deg(V)+\log N)).$
Note that right-hand side of the inequality in Theorem 10 depends very weakly on the embedding dimension $N$ ($\log N$ instead of $N$) compared to the bound in Theorem H. Notice also that while the bound in Theorem H is in terms of the degrees of the real polynomials cutting out $V$, the bound in Theorem 10 is in terms of the degree of $V$ (as a complex variety) which is a more intrinsic invariant. It is well known that any affine variety of degree $D$ in $\mathrm{C}^{N}$ can be cut out by at most $N+1$ polynomials of degree at most $D$. However, if we use this fact in conjunction with Theorem H, we will end up with a much weaker bound than the one in Theorem 10 ($N$ instead of $\log N$ in the expression on the right-hand side).
3.3. Lower bounds for algebraic computation trees
We briefly recall the definition of an algebraic computation tree (see for example [19] from which the following description is taken).
Definition 3.3.
An algebraic computation tree $T$ with input variables $X_{1},\ldots,X_{N}$ is a tree having three kinds of vertices:
  1. (a)
    computation vertices having outdegree $1$;
  2. (b)
    branching vertices with outdegree $3$, and the tree outgoing edges labelled by 00, $1$ and $-1$;
  3. (c)
    leaves with outdegree 00 and labelled by “YES” or “NO”.
[…]
At each branch vertex $v$, the variable $Y_{v}$ is assigned the value which is either a real constant, or an input variable, or a variable associated with a predecessor vertex. The three outgoing edges of a branch vertex $v$ labelled 00, $1$ or $-1$, then correspond to signs $Y_{v}=0$, $Y_{v}>0$, $Y_{v}<0$ (respectively).
With each leaf $w$, we associate naturally a a basic semi-algebraic set $S_{w}\subset\mathbb{R}^{N}$, defined by a conjunction of inequalities of the type $Y_{v}>0$, $Y_{v}=0$, $Y_{v}<0$, where $v$ varies over all branch nodes on the path from from the root to the leaf $w$ and the sign of $Y_{v}$ corresponds to the sign of the branch taken by the path. The semi-algebraic set $S_{T}$ associated to the tree $S$ is defined as the union
$S_{T}=\bigcup_{w\in\mathrm{Leaves}_{1}(T)}S_{w},$
where $\mathrm{Leaves}_{1}(T)$ denotes the set of leaf nodes of $T$ labelled “YES”.
We say that the algebraic computation tree $T$ tests membership in the semi-algebraic set $S$. The height of a tree is the length of the longest path from its root to a leaf node.
The interpretation of this model of computation is as follows. On an input
$x=(x_{1},\ldots,x_{N})\in\mathbb{R}^{N},$
the input variables $X_{i}$ get the corresponding values $x_{i}$, the arithmetic operations are executed in computation vertices $v$, and $Y_{v}$ is assigned the corresponding real value.
At branch vertices, the outward branch is chosen corresponding to the assigned value of $Y_{v}$. The input $x$ belongs to the semi-algebraic set $S_{w}$ if this process ends in the leaf $w$. If $w$ is a leaf marked “YES”, then $x$ is said to be accepted by the tree $T$.
It was proved by by Ben-Or [12] that for any semi-algebraic set the height of an algebraic computation tree testing membership in a semi-algebraic set $S\subset\mathrm{R}^{N}$, is at least $c_{1}\log b_{0}(S)-c_{2}N$, where and $c_{2},c_{2}>0$ absolute constants. Yao [50] proves that for any locally closed and bounded semi-algebraic set $S\subset\mathrm{R}^{N}$, the height of any algebraic computation tree testing membership in $S$ is bounded by
$c_{1}\log b^{BM}(S)-c_{2}N,$
where $b^{BM}(S)$ denotes the Borel-Moore Betti numbers of $S$ (they can also be defined as the dimensions of the compactly supported cohomology groups $\mbox{\rm H}_{c}^{*}(S)$). This result was extended recently by Gabrielov and Vorobjov [19] who proved that the height of an algebraic computation tree testing membership in a semi-algebraic set $S\subset\mathrm{R}^{N}$ (not necessarily locally closed), is at least
$c_{1}\frac{\log b_{m}(S)}{m+1}-c_{2}\frac{N}{m+1},$
where $c_{1},c_{2}>0$ are absolute constants.
We prove that the following theorem which in particular improves the original result of Ben-Or in a direction that is independent of [50] and [19].
Theorem 11.
Let $S\subset\mathrm{R}^{N}$ be a semi-algebraic set contained in an real algebraic set $V$. Suppose that $V$ is a union of non-singular real algebraic varieties, then the height of an algebraic computation tree testing membership in $S$, is at least
$\frac{1}{(1+(\log\deg V))}(c_{1}\log b_{0}(S)-c_{2}\dim V),$
where and $c_{1},c_{2}>0$ absolute constants.
Remark 14.
Note that the above lower bounds are independent of the ambient dimension $N$. As in previous remarks this is useful in situations where $\dim V$ and $\log\deg V$ are very small compared to the dimension of the ambient space.
Example 3.3.
For example, consider the Grassmannian variety $\mathrm{Gr}(m,N)$ of $m$-dimensional subspaces in $\mathrm{C}^{N}$, as a subvariety of $\mathbb{P}_{\mathrm{C}}^{\binom{N}{m}-1}$. Let $V=V_{m,N}\subset\mathrm{C}^{\binom{N}{m}}$ be the affine cone over $\mathrm{Gr}(m,N)$. Then, it is well known (see for example [33, Corollary 3.2.14]) that
(3.5) $\deg(V_{m,N})=\deg(\mathrm{Gr}_{m,N})=\frac{0!1!\cdots(n-1)!}{(m+1)!\cdots(m+n-1)!}(mn)!,$
where $n=N-m$, while
$\dim V_{m,N}=m(N-m)+1.$
This implies that $\deg(V_{m,N})=O(1)^{m(N-m)}$, and hence $\log(\deg(V_{m,N}))=O(m(N-m))$.
Now suppose that $S\subset V_{m,N}(\mathrm{R})$ is a semi-algebraic subset. Then, the lower bound in Theorem 11 yields that the depth of any algebraic computation tree testing membership in $S$ is at least
$c_{1}\frac{\log b_{0}(S)}{m(N-m)}-c_{2},$
while the result of Ben-Or implies a lower bound of
$c_{1}\log b_{0}(S)-c_{2}\binom{N}{m}$
(with different constants $c_{1},c_{2}$).
If $m>2$ and $b_{0}(S)<\frac{c_{2}}{c_{1}}\binom{N}{m}$, then the last inequality will not give a non-trivial lower bound, while Theorem 11 will yield a non-trivial bound.
3.4. Relative ranks in vector spaces and algebras
The notion of rank of tensors play a very important role in many applications – from understanding the complexity of matrix multiplication, to machine learning and quantum information theory (see for example [30] for an overview). We introduce here a notion of relative rank for finite dimensional vector spaces and algebras (relative to an arbitrary algebraic subset $\Delta$ of the vector space or algebra) which generalizes the notion of rank for tensors. We prove lower bounds on the relative rank for algebraically defined subsets $W$ of the vector space or algebra, in terms of the dimension and degree of $\Delta$ and the cardinality (or the number of connected components) of the set (Theorems 12, 13, 14 and 15 below).
Note that while proving upper bounds on the classical rank of specific tensors (such as the matrix multiplication tensor) [30] is an important and active area of research, proving lower bounds on the rank of sets of tensors (particularly in the relative setting) has not been equally well studied. We believe the results stated below will prove useful in applications in addition to quantum complexity theory that we discuss here.
3.4.1. Relative rank in vector spaces
Definition 3.4 (Relative rank in vector spaces).
Let $V$ be a $k$-vector space, and $\Delta\subset V$. For $v\in V$, we define
$\mathrm{rank}_{V,\Delta}(v)$
as the smallest number $t$, if it exists, such that there exists $v_{1},\ldots,v_{t}\in\Delta$ such that $v=v_{1}+\cdots+v_{t}$, and $\infty$ otherwise. We will call $\mathrm{rank}_{V,\Delta}(v)$ the rank of $v$ relative to $\Delta$.
For any subset $W\subset V$, we will denote
$\mathrm{rank}_{V,\Delta}(W)=\max_{w\in W}\mathrm{rank}_{\Delta}(w).$
Remark 15.
Note that if $\Delta\subset\Delta^{\prime}$, then $\mathrm{rank}_{V,\Delta}(v)\geq\mathrm{rank}_{V,\Delta^{\prime}}(v)$ for all $v\in V$, and $\mathrm{rank}_{V,\Delta}(W)\geq\mathrm{rank}_{V,\Delta^{\prime}}(W)$ for all subsets $W\subset V$.
Remark 16.
Notice that Definition 3.4 generalizes the standard notion of rank of a tensor. For example, if $V=V_{1}\otimes\cdots\otimes V_{m}$, and $\Delta=\{v_{1}\otimes\cdots\otimes v_{m}\mid v_{i}\in V_{i},1\leq i\leq m\}$ the subset of tensors of rank at most one, then for all $v\in V$, $\mathrm{rank}_{V,\Delta}(v)$ is the usual rank of $v$ as an element of the tensor product $V_{1}\otimes\cdots\otimes V_{m}$.
Theorem 12.
Let $k$ be an algebraically closed field and $V$ a finite dimensional vector space. Let $\Delta\subset V$ be an algebraic subset with $\dim(\Delta)=p$ and $\deg(\Delta)=D$, and $\mathcal{P}\subset k[V]_{\leq\delta}$ with $\mathrm{card}(\mathcal{P})=s$. Then, for any finite $\mathcal{P}$-constructible subset $W\subset A$
$\mathrm{rank}_{V,\Delta}(W)\in\Omega\left(\frac{\log(\mathrm{card}(W))}{p(\log\delta+\log s)+\log D}\right)$
(where the constant in the $\Omega(\cdot)$ is an absolute constant).
In the real closed case we have a slightly worse lower bound (with an added non-singularity assumption on $V$).
Theorem 13.
Let $\mathrm{R}$ be an real closed field and $V$ a finite dimensional $\mathrm{R}$-vector space. Let $\Delta\subset V$ be an algebraic set, which is a union of non-singular real varieties, with $\dim(\Delta)=p$ and $\deg(\Delta)=D$, and $\mathcal{P}\subset\mathrm{R}[V]_{\leq\delta}$ with $\mathrm{card}(\mathcal{P})=s$. Then, for any finite $\mathcal{P}$-semi-algebraic subset $W\subset V$
$\mathrm{rank}_{V,\Delta}(W)\in\Omega\left(\frac{\log(\mathrm{card}(W))}{p(\log\delta+\log s+\log D)}\right)$
(where the constant in the $\Omega(\cdot)$ is an absolute constant).
Remark 17.
Note that we cannot use Theorem 12 with $k=C=\mathrm{R}[i]$ to deduce Theorem 13 as a special case. A finite $\mathcal{P}$-semi-algebraic set need not be realizable as a finite $\mathcal{Q}$-constructible set over $\mathrm{C}$, where the cardinality and the degrees of the polynomials in $\mathcal{Q}$ are proportional to those of $\mathcal{P}$.
3.4.2. Relative rank in algebras
Let $k$ be a field and $A$ a finite-dimensional $k$-algebra. The algebra $A$ is allowed to be non-commutative or non-associative and we do not require $A$ to be unital. For example, $A$ could be a finite-dimensional Lie algebra, the algebra of endomorphisms of a finite dimensional $k$-vector space, the group algebra $k[G]$ for some finite group $G$ (or any sub-algebra of $k[G]$). In each of these cases, it is meaningful to ask given some subset $\Delta$ of $A$ and an element $f\in A$, how many elements of $\Delta$ does one need to express $f$ as a product of elements of $\Delta$.
Definition 3.5 (Relative rank in algebras).
Let $A$ be a $k$-algebra and $\Delta\subset A$. For $\alpha\in A$, we define
$\mathrm{rank}_{A,\Delta}(\alpha)$
as the smallest number $t$, if it exists, such that there exists $\alpha_{1},\ldots,\alpha_{t}\in\Delta$ such that $\alpha=\alpha_{1}\cdots\alpha_{t}$, and $\infty$ otherwise. We will call $\mathrm{rank}_{A,\Delta}(\alpha)$ the rank of $\alpha$ relative to $\Delta$.
For any subset $W\subset A$, we will denote
$\mathrm{rank}_{A,\Delta}(W)=\max_{\alpha\in W}\mathrm{rank}_{A,\Delta}(\alpha).$
Remark 18.
Note that if $\Delta\subset\Delta^{\prime}$, then $\mathrm{rank}_{A,\Delta}(\alpha)\geq\mathrm{rank}_{A,\Delta^{\prime}}(\alpha)$ for all $\alpha\in A$, and $\mathrm{rank}_{A,\Delta}(W)\geq\mathrm{rank}_{A,\Delta^{\prime}}(W)$ for all subsets $W\subset A$.
Definition 3.6 (Relative max-min rank in real algebras).
More generally, let $W$ be a semi-algebraic subset of $A$ (not necessarily finite). We define the max-min-rank, $\mathrm{mmrank}_{A,\Delta}(W)$ of $W$ relative to $\Delta$ by
$\mathrm{mmrank}_{A,\Delta}(W)=\max_{C\in\mathrm{Cc}(W)}\min_{\alpha\in C}\mathrm{rank}_{A,\Delta}(\alpha).$
(Note that if $W$ is finite $\mathrm{mmrank}_{A,\Delta}(W)=\mathrm{rank}_{A,\Delta}(W)$.)
Notice that a finite dimensional $k$-algebra $A$ has the structure of a vector space (by forgetting the multiplication) and it makes sense to consider algebraic subsets of $A$. We denote the ring of polynomials on $A$ by $k[A]$ and the subspace of polynomials in $k[A]$ of degree at most $\delta$ by $k[A]_{\leq\delta}$.
Theorem 14.
Let $k$ be an algebraically closed field and $A$ a finite dimensional $k$-algebra. Let $\Delta\subset A$ be an algebraic subset with $\dim(\Delta)=p$ and $\deg(\Delta)=D$, and $\mathcal{P}\subset k[A]_{\leq\delta}$ with $\mathrm{card}(\mathcal{P})=s$. Then, for any finite $\mathcal{P}$-constructible subset $W\subset A$
$\mathrm{rank}_{A,\Delta}(W)\in\Omega\left(\frac{\log(\mathrm{card}(W))}{p(\log\delta+\log s+\log\log\mathrm{card}(W))+\log D}\right)$
(where the constant in the $\Omega(\cdot)$ is an absolute constant).
We have the following theorem in the real closed case.
Theorem 15.
Let $\mathrm{R}$ be a real closed field and $A$ a finite dimensional $\mathrm{R}$-algebra. Let $\Delta\subset A$ be an algebraic subset, with $\dim\Delta=p$, $\deg\Delta=D$, and suppose that $\Delta$ is a finite union of non-singular complete intersection real algebraic varieties. Let $\mathcal{P}\subset\mathrm{R}[A]_{\leq\delta}$ with $\mathrm{card}(\mathcal{P})=s$. Let $W\subset A$ be a $\mathcal{P}$-semi-algebraic subset. Then,
$\mathrm{mmrank}_{A,\Delta}(W)\in\Omega\left(\frac{\log(b_{0}(W))}{p(\log\delta+\log s+\log\log b_{0}(W)+\log D)}\right)$
(where the constant in the $\Omega(\cdot)$ is an absolute constant).
3.4.3. Example application of Theorem 12
Example 3.4 (Ranks of tensors).
Let $k$ be algebraically closed and let $V_{1},\ldots,V_{m}$ be finite dimensional vector spaces with $\dim V_{i}=N_{i}$. Let $V=V_{1}\otimes\cdots\otimes V_{m}$. For $1\leq i\leq m$, let $B_{i}=(e_{i,j})_{1\leq j\leq N_{i}}$ be a basis of $V_{i}$, and $B=B_{1}\otimes\cdots\otimes B_{m}$ the induced basis of $V$. The coordinates of an element $v$ with respect to the basis $B$ is then given by a $m$-dimensional array in $k^{N_{1}\times\cdots\times N_{m}}$. Let $B^{\prime}\subset B$ be a fixed set and $W\subset V$, be defined as the finite subset of elements of $V$, whose coordinates with respect to the basis elements in $B^{\prime}$ belong to $\{0,1\}$ and whose coordinates with respect to the basis elements $B-B^{\prime}$ are 00. Clearly, $\mathrm{card}(W)=2^{\mathrm{card}(B^{\prime})}$. (In terms of arrays $W$ is a finite set of $N_{1}\times\cdots\times N_{m}$ arrays whose entries in the positions indexed by $B^{\prime}$ are 00 or $1$ and whose remaining entries are 00).
Let $\Delta\subset V$ be the set of tensors of rank $\leq 1$. Then, $\Delta$ can be identified with the affine cone over the image of the Segre embedding
$\mathbb{P}_{k}^{N_{1}-1}\times\cdots\mathbb{P}_{k}^{N_{m}-1}\rightarrow\mathbb{P}_{k}^{N_{1}\cdots N_{m}-1},$
which yields:
$p=\dim\Delta=N_{1}+\cdots+N_{m}-m+1,$
and
$D=\deg\Delta=\binom{N_{1}+\cdots+N_{m}-m}{N_{1}-1,\ldots,N_{m}-1}$
using the multinomial notation (see for example [16, Proposition 2.11]). Applying Theorem 12, and noting that $\mathrm{card}(W)=2^{\mathrm{card}(B^{\prime})}$, we obtain
$\mathrm{rank}_{V,\Delta}(W)=\Omega\left(\frac{\mathrm{card}(B^{\prime})}{p\cdot(\log(N_{1}\cdots N_{m})+\log D)}\right).$
Taking $m=2$, and $N_{1}=N_{2}=N$, and $B^{\prime}\subset B$ with $\mathrm{card}(B^{\prime})=N^{1+{\varepsilon}}$, for some ${\varepsilon},0<{\varepsilon}<1$, we obtain that
$\mathrm{rank}_{V,\Delta}(W)=\Omega\left(\frac{N^{{\varepsilon}}}{\log N}\right).$
Thus, in any set of $N\times N$ zero-nonzero matrices with non-zero entries allowed only in a fixed set of $N^{1+{\varepsilon}}$ positions, there exists a matrix of rank at least $C\cdot\frac{N^{\varepsilon}}{\log N}$, where $C>0$ is an absolute constant.
Note that a possible alternative approach to proving the above statement is to use the well known Kővári-Sós-Turán (KST) theorem (see for example [52, Theorem 1.4.2]). Identifying the set $B$ with $[1,N]\times[1,N]$, if the chosen set of positions, $B^{\prime}\subset B$ contains a “combinatorial square” of size $s\times s$ (i.e. $I\times J$, with $\mathrm{card}(I)=\mathrm{card}(J)=s,I,J\subset[1,N]$), then clearly, $\mathrm{rank}_{V,\Delta}(W)\geq s$. The KST theorem guarantees the existence of such a square if $\mathrm{card}(B^{\prime})\geq C(s)\cdot N^{2-1/s}$. In the KST theorem $s$ is constant. For $s$ growing with $N$, the cardinality of $B^{\prime}$ needs to be even larger. Thus, for $s=N^{\varepsilon}/\log N$, $\mathrm{card}(B^{\prime})$ needs to be at least $N^{2-o(1)}$ to be guaranteed that $B^{\prime}$ contains a combinatorial square with $s=N^{\varepsilon}$. However, $N^{1+{\varepsilon}}\ll N^{2-o(1)}$, and thus this approach cannot be used to prove the claim above.
3.4.4. Example of application of Theorem 14
Example 3.5 (Lengths of permutations).
Let $V$ be an $n$-dimensional $k$-vector space and $A=\mathrm{End}(V)$. Let $B=\{v_{1},\ldots,v_{n}\}$ be a vector space basis of $A$, and $W$ the subset of $A$ of those endomorphism that permutes the elements of $B$. Then, $W$ is a zero-dimensional algebraic subset of $A$ cut out by $n^{2}+2n$ equations of degrees at most $2$ ($n^{2}$ equations of degree $2$ and $2n$ equations of degree $1$), and $\mathrm{card}(W)=n!$.
If $\Delta\subset A$ be any $p$-dimensional variety having degree $D$, then using Theorem 1 we have:
$\mathrm{rank}_{A,\Delta}(W)=\Omega\left(\frac{n\log n}{(p+1)\log n+\log D}\right).$
This bound is asymptotically tight.
Consider for example $A=\mathrm{End}(V)$, with $V$ a $k$-vector space with basis $v_{1},\ldots,v_{n}$, and $\Delta$ the union of $\mathrm{Id}_{V}$, and the endomorphisms $\pi_{i},1\leq i\leq n-1$, where $\pi_{i}$ acts on $B$ by transposing $v_{i},v_{i+1}$ and by the identity on $v_{j},j\not\in\{i,i+1\}$. Then as is well known each element of $W$ can be expressed as a product of at most $2N$ elements of $\mathrm{Id}_{V},\pi_{1},\ldots,\pi_{N-1}$.
So, $\mathrm{rank}_{A,\Delta}(W)=O(n)$, while in the lower bound above we have $p=0$, $D=n$, which gives $\mathrm{rank}_{A,\Delta}(W)=\Omega(n)$. Indeed the $\Omega(n)$ lower bound is valid for $p=O(1)$ and $D\leq n$.
3.5. Quantum circuit lower bounds
3.5.1. Shannon’s lower bound for classical circuits
Boolean circuit complexity measures complexity of Boolean functions $f:\{0,1\}^{n}\rightarrow\{0,1\}$ by the size of the smallest circuit computing $f$. We assume that the gates of a circuit come from some fixed finite universal set of gates. We refer the reader to [49, Page 7, Definition 3.1] for the precise definition of a circuit. Notice that in this case the set of circuits of bounded size is finite.
Shannon [40] showed using a counting argument (using the fact that as noted before that the set of circuits of bounded size is finite) very early that almost all Boolean functions need circuits of size $\Omega({2^{n}}/{n})$ (see [49, Page 90, Theorem 2.1]. Here almost all refers to the fact that the number of Boolean functions that require circuits of size $\Omega({2^{n}}/{n})$ is bounded from below by $2^{2^{n}}(1-o(1))$. It is also known (see for instance [49, Page 92, Theorem 2.2]) that every Boolean function $f:\{0,1\}^{n}\rightarrow\{0,1\}$ can be computed by a circuit of size $O({2^{n}}/{n})$.
Remark 19.
In the theorems cited above the set of gates used is assumed to be all possible gates of arity two (i.e. the finite set of $162^{2^{2}}=$ 16 gates computing all possible Boolean functions $g:\{0,1\}^{2}\rightarrow\{0,1\}$). The arity two is not important. The same asymptotic results will hold even if we allow all gates of some bounded arity $q\geq 2$ (the constants will depend on the arity) .
3.5.2. Analog of Shannon’s lower bound for quantum circuits
We begin with a brief review of quantum complexity since it motivates our definition. Recall that in quantum complexity theory a qubit refers to a two-dimensional complex Hermitian vector space $V$ with a fixed orthonormal basis =$\mathcal{B}=\{|0{\rangle},|1{\rangle}\}$. A $n$-qubit register refers to the $n$-fold tensor power, $V^{\otimes n}$ of $V$, with an induced Hermitian structure in which the vectors $\mathcal{B}^{\otimes n}=\{|j{\rangle}|0\leq j\leq N-1\}$ form an orthonormal basis (here $N=2^{n}$, and we identify a 00-$1$ string $x_{n-1}\cdots x_{0}$ of length $n$, with the integer
$x=\sum_{k=0}^{n-1}x_{k}\cdot 2^{k},$
and denote $|x_{n-1}{\rangle}\otimes\cdots\otimes|x_{0}{\rangle}$ simply by $|\mathbf{x}{\rangle}$).
A $q$-ary quantum gate $g$ is an element $U_{g}\in\mathbf{U}(V^{\otimes q})$. In a quantum circuit $C$, with input $n$-qubits, such a gate operates on a subset of $q$ of the $n$-qubits (say labelled by the tuple $\mathbf{i}=(i_{1},\ldots,i_{q}),1\leq i_{1}<\cdots<i_{q}\leq$), which gives rise to an element $U_{g,\mathbf{i}}$ of $\mathbf{U}(V^{\otimes n})$ which can be identified with $U_{g}\otimes\mathrm{Id}_{V^{\otimes(n-q)}}$. (after rearranging the order of the copies of $V$ in the tensor power $V^{\otimes n}$ bringing the ones indexed by $(i_{1},\ldots,i_{q})$ to the the positions $(1,\ldots,q)$).
A quantum circuit $C$ with $n$ input qubits labelled by $1,\ldots,n$ having $r$ $q$-ary gates is determined by the following data:
  1. 1.
    an ordering of the $r$ gates (lets suppose the ordered tuple of gates is $g_{1},\ldots,g_{r}$, and
  2. 2.
    for each $i,1\leq i\leq r$, an ordered choice, $\mathbf{i}=(i_{1},\ldots,i_{q})$ of $q$ elements from amongst $1,\ldots,n$.
We denote
$U(C)=\prod_{1\leq i\leq r}U_{g_{i},\mathbf{i}}\in\mathbf{U}(V^{\otimes n})$
the unitary transformation implemented by $C$ (where $N=2^{n}$), and we denote $\mathrm{size}(C)=r$.
For $q\geq 1$, we will denote by $\mathcal{C}_{q}$, the set of quantum circuits using $q$-ary quantum gates.
Definition 3.7 (Quantum circuit complexity of a unitary transformation).
The $q$-ary quantum circuit complexity of a unitary transformation $\mathbf{U}(V^{\otimes n})$ is defined as
$\min_{C\in\mathcal{C}_{q},U(C)=U}\mathrm{size}(C).$
In other words, if we denote
$\Delta=\Delta_{n,q}=\{U_{g,\mathbf{i}}\mid U_{g}\in\mathbf{U}(V^{\otimes q}),\mathbf{i}=(i_{1},\ldots,i_{q}),1\leq i_{1}<\cdots<i_{q}\},$
then the $q$-ary quantum circuit complexity of a unitary transformation $U\in\mathbf{U}(V^{\otimes n})$ is precisely (using Definition 3.5)
$\mathrm{rank}_{\mathrm{End}(V^{\otimes n}),\Delta}(U)$
considering $U$ as an element of $\mathrm{End}(V^{\otimes n})$. This motivates the following definition:
Definition 3.8 (Relative quantum circuit complexity).
For arbitrary $\Delta\subset\mathrm{End}(V^{\otimes n})$ we will call $\mathrm{rank}_{\mathrm{End}(V^{\otimes n}),\Delta}(U)$ the quantum complexity of $U$ relative to $\Delta$.
Remark 20.
In analogy with classical complexity, $\Delta$ could be thought of as a set of quantum oracles – with unit cost for each application of an element of $\Delta$ in an quantum circuit. Notice though we do not assume that $\Delta$ is contained in the unitary group $\mathbf{U}(V^{\otimes n})$, but it is allowed to be.
We now explain (following [32]) what is meant by a quantum circuit computing a Boolean function. For ease of understanding, we start with a provisional (very stringent) definition and then give a more general (much less stringent) definition afterwards.
Let $f:\{0,1\}^{n}\rightarrow\{0,1\}$ be a Boolean function. We associate a unitary transformation $U_{f}\in\mathbf{U}(2^{n+1})$ to $f$ which takes for $x_{0},\ldots,x_{n-1},y\in\{0,1\}$, the separable state $|\mathbf{x}{\rangle}\otimes|y{\rangle}$ to the state $|\mathbf{x}{\rangle}\otimes|f(x)\oplus y{\rangle}$, where $\oplus$ denotes “exclusive-or”. Note that using the measurement postulate of quantum mechanics, if we set the input qubits to $|\mathbf{x}{\rangle}\otimes|0{\rangle}$ and measure the ancillary bit in the output, we will be left in the state $|1{\rangle}$ if $f(x)=1$ and in the state $|0{\rangle}$ if $f(x)=0$ with probability $1$.
Definition 3.9.
We will call the quantum circuit complexity of $U_{f}$ as the stringent quantum circuit complexity of $f$, and the quantum complexity of $U_{f}$ relative to $\Delta$ as stringent quantum complexity of $f$ relative to $\Delta$.
We prove the following theorem.
Theorem 16.
There exists $C>0$, such that for all $n>0$, $\Delta\subset\mathrm{End}(V^{\otimes(n+1)})$ algebraic subset with $\dim\Delta=p$ and $\deg\Delta=D$, there exists Boolean functions $f:\{0,1\}^{n}\rightarrow\{0,1\}$, such that the stringent quantum complexity of $f$ relative to $\Delta$, is at least
$C\cdot\frac{2^{n}}{p\cdot n+\log D}.$
We can deduce as a corollary (setting $p=2^{q}\times 2^{q})$ and $D=\binom{n}{q}$ in Theorem 16) the following lower bound which is the quantum analog of Shannon’s lower bound for classical circuits.
Corollary 1.
For each $q\geq 1$, there exists $C=C_{q},0<c<1$ (depending only on $q$), such that for each $n>0$ there exists Boolean functions $f:\{0,1\}^{n}\rightarrow\{0,1\}$, such that the stringent quantum circuit complexity of $f$ is least $C\cdot\frac{2^{n}}{n}$.
Remark 21.
It follows that the fraction of Boolean functions that need quantum circuits in $\mathcal{C}_{q}$ of size greater than $c\cdot\frac{2^{n}}{n}$ is $1-2^{-2^{n-1}}=1-o(1)$.
Remark 22.
Notice that since the set of gates $\mathcal{U}_{q}$ is infinite (uncountably so), and hence the number of distinct quantum circuits of any bounded size is also uncountably infinite. So a priori there is no reason for such circuits not being able to compute all Boolean functions on $n$ variables.
As mentioned before we will work with a more general notion of what it means for a quantum circuit to compute a Boolean function. We relax our prior definition in two ways. First, instead of ending up in the right state depending on the value of the function $f$ with probability $1$, we are satisfied if we end at the right state with probability $>1/2$. We will also allow more than one ancillary bits. The precise definition is as follows.
Definition 3.10 (Unitary transformation associated to a Boolean function).
We denote the $n$ input qubits by $x_{0},\ldots,x_{n-1}$ and $t+1$ ancillary qubits denoted by $y,z_{1},\ldots,z_{t}$.
Given a Boolean function $f:\{0,1\}^{n}\rightarrow\{0,1\}$, we denote by $U_{f,t}^{+}$ the set of all unitary transfomations $U$ of $V^{\otimes(n+1+1)}$ satisfying for each $x,x^{\prime},0\leq x,x^{\prime}\leq 2^{n}-1$, $y\in\{0,1\}$,
[…]
Remark 23.
Note that $U_{f}\in U_{f,1}^{+}$. Also note that the inequality appearing in (3.6) is the most relaxed possible in as much as we do not insist on any positive gap between the accepting and rejecting probabilities.
Definition 3.11.
We will call the minimum of the quantum circuit complexities of $U\in U_{f,t}^{+}$ as the quantum circuit complexity of $f$ with $t+1$ ancillary bits, and the minimum of quantum complexities of $U\in U_{f,t}^{+}$ relative to $\Delta$ as the quantum complexity of $f$ with $t+1$ ancillary bits relative to $\Delta$.
In the following theorem we consider $A=\mathrm{End}(V^{\otimes(n+t+1)})$ as a real subspace of the $4^{n+t+1}\times 4^{n+t+1}$ dimensional real vector space $A_{\mathbb{R}}=\mathrm{End}(V_{\mathbb{R}}^{\otimes(n+1+1)})$ considering $V$ as a $4$-dimensional $\mathbb{R}$-vector space denoted $V_{\mathbb{R}}$.
Theorem 17.
There exists $C>0$, such that for all $n>0$, $\Delta\subset A$, such that $\Delta$ is a real algebraic subset of $A_{\mathbb{R}}$, with $\dim\Delta=p$ and $\deg\Delta=D$, and suppose $\Delta$ is a union of real non-singular complete intersection varieties. Then, there exists Boolean functions $f:\{0,1\}^{n}\rightarrow\{0,1\}$, such that the quantum complexity of $f$ with $t+1$ ancillary bits relative to $\Delta$, is at least
$C\cdot\frac{2^{n}}{p\cdot(n+t+\log D)}.$
The following corollary of Theorem 17 (without the oracle $\Delta)$ already appeared in [37], and we include it here for completeness.
Corollary 2.
[37, Theorem 7.1] For each $q\geq 1$, there exists $C=C_{q},0<c<1$ (depending only on $q$), such that for each $n>0$ there exists Boolean functions $f:\{0,1\}^{n}\rightarrow\{0,1\}$, such that the quantum circuit complexity of $f$ with $t+1$ ancillary bits is at least $C\cdot\frac{2^{n}}{n+t}$.
Remark 24.
A more traditional approach would be to consider circuits built out of only Clifford and Toffoli gates (using results form [28, 39]) that approximate the required unitary matrix sufficiently well. Since the number of circuits of bounded size built out of a finite set of quantum gates is finite one can then hope to reproduce the original counting argument of Shannon in the classical case. This fails because of the following reason.
Our very generous definition of what it means for quantum circuit to compute a Boolean function gives no room for any approximation, as any error could mean that the new circuit $C^{\prime}$ does not compute $f$. If we strengthen the definition (thus make proving lower bounds easier) by insisting that there be a gap $\delta$ between the probabilities of acceptance/rejection i.e. we replace the inequality in Definition 3.10 by
[…]
$<$ $1/2-\delta\text{ otherwise}$
for some fixed $\delta,0\leq\delta<1/2$, then one can carry through the program sketched earlier using approximation and counting (see for example [29, 28]). In this case one would need to replace $C$ by a circuit $C^{\prime}$ using Clifford and Toffoli gates which approximates $U(C)$ within $\delta/2$ in max-norm. This produces the result that almost all Boolean functions need quantum circuits of size
$\Omega(\min(2^{n}/(n\log n),2^{n}/((\log(1/\delta)\log n)))).$
This lower bound goes to 00 as $\delta\rightarrow 0$. So it does not produce any meaningful lower bound for $\delta=0$ (which is the case in Definition 3.10). We also note that there is recent work in constructing explicit Boolean functions with exponential quantum circuit complexity using sophisticated tools from transcendental number theory [26]. Our rank argument does not produce such functions.
4. Proofs
4.1. Proof of Theorem 1
In our proof of Theorem 1 we will need to use the following generalized form of Bezout’s theorem (see for instance [18, Example 8.4.6, page 148]).
Theorem I (Generalized Bezout).
Let $V_{1},\ldots,V_{s}$ be algebraic subsets of some projective space over $k$, such that each $V_{i}$ is a union of varieties of the same dimension, and $Z_{1},\ldots,Z_{t}$ be the irreducible components of the intersection $V_{1}\cap\cdots\cap V_{s}$. Then
[…]
4.1.1. Outline of the proof Theorem 1
The main idea (which is very similar to that used in [25], and also in [8]) is to first bound the number of zero-non-zero patterns whose realizations are zero dimensional using Theorem I. The main subtlety here is that the dimension of an algebraic set need not decrease upon intersection with another hypersurface. The problem of bounding the number of zero-nonzero patterns whose realizations have dimension $>0$ then can be reduced to the zero-dimensional case by taking generic hyperplane sections.
The proof of Theorem 1 will now follow from the following two lemmas. We follow closely the proof of Lemma 7.3 in [8]. Observe that it is sufficient to prove the theorem in case $V$ is irreducible of dimension $p$. Otherwise, we can take the irreducible decomposition of $V$ noting that their degrees add up to $D$.
Notation 4.1.
For a finite set of polynomials $\mathcal{Q}\subset k[X_{1},\ldots,X_{N}]$, and an algebraic set $W\subset k^{N}$, we denote by ${\rm Z}(\mathcal{Q},W)$ the set of common zeros of $\mathcal{Q}$ in $V$. If $\mathcal{Q}=\{Q\}$, we write ${\rm Z}(Q,W)$ for ${\rm Z}(\mathcal{Q},W)$.
Let $V,\mathcal{P}$ be as in Theorem 1. For $\rho\in\{0,1\}^{\mathcal{P}}$, let ${\mathcal{P}}_{\rho}=\{P\in{\mathcal{P}}\mid\rho(P)=0\}$ and let $V_{\rho}={\rm Z}({\mathcal{P}}_{\rho},V)$. Let $V_{\rho}=V_{\rho,1}\cup\cdots\cup V_{\rho,n(\rho)}$ denote the decomposition of $V_{\rho}$ into irreducible components.
We first prove a bound on the number of isolated points occurring as ${\mathcal{R}}(\rho,V)$ for some $\rho\in\{0,1\}^{\mathcal{P}}$.
Lemma 4.1.
$\mathrm{card}(\{(\rho,i)\mid\dim V_{\rho,i}=0\})\leq\binom{s}{p}\cdot D\cdot d^{p}.$
Proof.
Let ${\mathcal{P}}=\{P_{1},\ldots,P_{s}\}$ (the implicit ordering of the polynomials in ${\mathcal{P}}$ will play a role in the proof).
First consider a sequence of $p$ polynomials of ${\mathcal{P}}$, $P_{i_{1}},\ldots,P_{i_{p}},$ with $1\leq i_{1}<\cdots<i_{p}\leq s$. Let $\{W_{\alpha_{1}}\}_{\alpha_{1}=1,2..}$ denote the irreducible components of ${\rm Z}(P_{i_{1}},V)$. Similarly, let $\{W_{\alpha_{1},\alpha_{2}}\}_{\alpha_{2}=1,2,\ldots}$ denote the irreducible components of $W_{\alpha_{1}}\cap{\rm Z}(P_{i_{2}},V)$ and so on. We prove by induction that for $1\leq\ell\leq p$
(4.1) $\sum_{\alpha_{1},\alpha_{2},\ldots,\alpha_{\ell}}\deg(W_{\alpha_{1},\ldots,\alpha_{\ell}})\leq D\cdot d^{\ell}.$
The claim is trivially true for $\ell=1$. Suppose it is true up to $\ell-1$. Then using Theorem I we get that
$\sum_{\alpha_{1},\ldots,\alpha_{\ell-1}}\deg({\rm Z}(P_{\ell},W_{\alpha_{1},\ldots,\alpha_{\ell-1}}))\leq D\cdot d^{\ell-1}\cdot d=D\cdot d^{\ell}.$
Now suppose that $\dim V_{\rho,i}=0$ and let $V_{\rho,i}=\{x\}$.
We first claim that there must exist $P_{i_{1}},\ldots,P_{i_{p}}\in{\mathcal{P}}_{\rho}$, with $1\leq i_{1}<i_{2}<\cdots<i_{p}\leq s$, and irreducible affine algebraic varieties $V_{0},V_{1},\ldots,V_{p}$ satisfying:
  1. (A)
    $V=V_{0}\supset V_{1}\supset V_{2}\supset\cdots\supset V_{p}=\{x\}$, and
  2. (B)
[…]
Note that in the sequence, $V_{0},V_{1},\ldots,V_{p}$, $\dim(V_{i})$ is necessarily equal to $p-i$. To prove the claim, let
${\mathcal{P}}_{\rho}=\{P_{j_{1}},\ldots,P_{j_{m}}\},$
with $1\leq j_{1}<j_{2}<\cdots<j_{m}\leq s$. Let $W_{0}=V$ and for each $h,1\leq h\leq m$ define inductively $W_{h}$ to be an irreducible component of ${\rm Z}(P_{j_{h}},W_{h-1})$ containing the point $p$. By definition, $V=W_{0}\supset W_{1}\supset\cdots\supset W_{m}=\{p\}$ and each $W_{h}$ is irreducible. Hence, there must exist precisely $p$ indices, $1\leq\alpha_{1}<\cdots<\alpha_{p}\leq m$ such that $\dim W_{\alpha_{i}}=\dim W_{\alpha_{i-1}}-1$. If $\beta\not\in\{\alpha_{1},\ldots,\alpha_{p}\}$, then $W_{\beta}=W_{\beta-1}$. Now let $i_{1}=j_{\alpha_{1}},\ldots,i_{p}=j_{\alpha_{p}}$, and $V_{j}=W_{i_{j}}$ for $0\leq j\leq p$. It is clear from construction that the sequence $V_{0},\ldots,V_{p}$ satisfies the properties stated above, thus proving the claim.
The number of sequences of polynomials of length $p$, $P_{i_{1}},\ldots,P_{i_{p}},$ with $1\leq i_{1}<\cdots<i_{p}\leq s$, is $\binom{s}{p}$. The lemma now follows from the bound in (4.1). ∎
The following lemma is a generalization of the above lemma to positive dimensions.
Lemma 4.2.
For $0\leq\ell\leq p$
[…]
Proof.
In order to bound the degree of $V_{\rho,i}$ with $\dim V_{\rho,i}=\ell$ it suffices to count the isolated zeros of the intersection of $V_{\rho,i}$ with a generic affine subspace $L$ of dimension $p-\ell$. Since there are only a finite number of $V_{\rho,i}$’s to consider we can assume that
  1. (A)
    $L$ is generic for all of them,
  2. (B)
    the sets $V_{\rho,i}\cap L$ are pairwise disjoint,
  3. (C)
    for each $V_{\rho,i}$ with $\dim_{\mathrm{C}}V_{\rho,i}=\ell$, ${\rm Z}(P,L\cap V_{\rho,i})=\emptyset$ for each $P\in{\mathcal{P}}\setminus{\mathcal{P}}_{\rho}$.
Now restrict to the subspace $L$ and apply Lemma 4.1, noting that an isolated point of $V_{\rho,i}\cap L$ is an isolated point of the family ${\mathcal{P}}$ restricted to $L$. ∎
Proof of Theorem 1.
Follows directly from Lemma 4.2. ∎
4.2. Proofs of Theorems 2, 3 and 4
We first recall some preliminary definitions and results.
4.2.1. Multiplicity and singular points of a variety
Given a variety $V$, and a point $x\in V$, the multiplicity of $V$ at $x$, denoted $\mathrm{mult}_{x}(V)$, is defined as the degree of the exceptional divisor of the blowup $\widetilde{V}$ of $V$ at $x$ [18, Section 4.3, page 79]. This definition as equivalent to the definition of the Samuel multiplicity of the local ring $\mathcal{O}_{V,x}$ [18, Example 4.3.1].
For an algebraic set $V$ with irreducible components $V_{1},\ldots,V_{m}$, and for $x\in V$, we define $\mathrm{mult}_{x}(V)=\sum_{i=1}^{m}\mathrm{mult}_{x}(V_{i})$ (where we define $\mathrm{mult}_{x}(V_{i})=0$ if $x\not\in V_{i}$).
Given a variety $V$, and a point $x\in V$, $\mathrm{mult}_{x}(V)=1$ if and only if the local ring $\mathcal{O}_{V,x}$ is regular [18, Example 4.3.5 (d)], and this is equivalent to $V$ being non-singular at $x$ [18, Section B.2.7, Page 429].
The following lemma follows immediately from the above which we record for future use.
Lemma 4.3.
If $V$ is an algebraic set, and $x\in V$. Then, $x$ is a non-singular point of $V$ if and only if $\mathrm{mult}_{x}(V)=1$.
The set of points of $V$ which are not non-singular will be denoted by $\mathrm{Sing}(V)$. It is a proper closed subset of $V$ (see [24, Theorem 5.3]). If $V$ is a variety, then $\dim\mathrm{Sing}(V)<\dim V$.
We say that $V$ is non-singular if $\mathrm{Sing}(V)=\emptyset$.
We also recall below the notion of an extension of a real closed field $\mathrm{R}$ by the real closed field $\mathrm{R}{\langle}\zeta{\rangle}$ of algebraic Puiseux series in $\zeta$ (where $\zeta$ is an infinitesimal i.e. $\zeta$ is positive but smaller than all positive elements of $\mathrm{R}$). These extensions are used for making infinitesimal perturbations of the given polynomials.
4.2.2. Real closed extensions and Puiseux series
We will need some properties of Puiseux series with coefficients in a real closed field. We refer the reader to [7] for further details.
Notation 4.2.
For $\mathrm{R}$ a real closed field we denote by $\mathrm{R}\left\langle{\varepsilon}\right\rangle$ the real closed field of algebraic Puiseux series in ${\varepsilon}$ with coefficients in $\mathrm{R}$. We use the notation $\mathrm{R}\left\langle{\varepsilon}_{1},\ldots,{\varepsilon}_{m}\right\rangle$ to denote the real closed field $\mathrm{R}\left\langle{\varepsilon}_{1}\right\rangle\left\langle{\varepsilon}_{2}\right\rangle\cdots\left\langle{\varepsilon}_{m}\right\rangle$. Note that in the unique ordering of the field $\mathrm{R}\left\langle{\varepsilon}_{1},\ldots,{\varepsilon}_{m}\right\rangle$, $0<{\varepsilon}_{m}\ll{\varepsilon}_{m-1}\ll\cdots\ll{\varepsilon}_{1}\ll 1$.
Notation 4.3.
For elements $x\in\mathrm{R}\left\langle{\varepsilon}\right\rangle$ which are bounded over $\mathrm{R}$ we denote by $\lim_{{\varepsilon}}x$ to be the image in $\mathrm{R}$ under the usual map that sets ${\varepsilon}$ to 00 in the Puiseux series $x$.
Notation 4.4.
If $\mathrm{R}^{\prime}$ is a real closed extension of a real closed field $\mathrm{R}$, and $S\subset\mathrm{R}^{k}$ is a semi-algebraic set defined by a first-order formula with coefficients in $\mathrm{R}$, then we will denote by ${\rm Ext}(S,\mathrm{R}^{\prime})\subset\mathrm{R}^{\prime k}$ the semi-algebraic subset of $\mathrm{R}^{\prime k}$ defined by the same formula. It is well known that ${\rm Ext}(S,\mathrm{R}^{\prime})$ does not depend on the choice of the formula defining $S$ [7, Proposition 2.87].
4.2.3. Outline of the proof of Theorem 2
We first observe that it suffices to consider the case when $V$ is irreducible of degree $D$ and dimension $p$, satisfying that $\dim_{x}V(\mathrm{R})=p$ for all $x\in V(\mathrm{R})$. We introduce an infinitesimal $\gamma_{0}$. Using the conic structure theorem for semi-algebraic sets [7, Proposition 5.49] we are able to restrict attention to realizations of sign conditions which are contained in a ball of radius $(1/\gamma_{0})^{1/2}$. This necessitates introduction of a new polynomial denoted by $P_{0}$ in the proof. We then define infinitesimal perturbations (using $s+2$ additional infinitesimals $\gamma_{1},\ldots,\gamma_{s},{\varepsilon},\delta$) of the polynomials in $\mathcal{P}$ and prove that it suffices to bound the number of semi-algebraically connected components of the realizations on $\widetilde{S}:={\rm Ext}(V(\mathrm{R}),\widetilde{\mathrm{R}})$ (the extension of $V(\mathrm{R})$ to the field $\widetilde{\mathrm{R}}$ obtained from $\mathrm{R}$ by adjoining infinitesimal elements) of a certain set of weak sign conditions on this new set of $14s+$ 1 polynomials (Lemma 4.4). We then show that in order to bound these, it suffices to bound the number of semi-algebraically connected components of all the non-empty real algebraic sets defined by the perturbed set of $14s+$ 1 polynomials which are contained in $\widetilde{S}$ (Lemma 4.7).
If we consider the real algebraic set defined as the intersection of $\widetilde{S}$ with $i$ of the perturbed polynomials, then we get a real algebraic set of dimension $p-i$ of degree bounded by $Dd^{i}$. So it now suffices to bound the number of semi-algebraically connected components of each of these real algebraic sets. Taking a sum of squares we reduce to the case where the real algebraic set (say) $\widetilde{Z}$, is defined as the intersection with $\widetilde{S}$ with the ${\rm Z}(\widetilde{Q},\widetilde{\mathrm{R}}^{N})$ for a single polynomial $\widetilde{Q}$.
We bound $b_{0}(\widetilde{Z})$ in two steps.
  1. (a)
    We first prove (using the semi-algebraic curve selection lemma and the intermediate value theorem for real closed fields) that
    $b_{0}(\widetilde{Z})\leq b_{0}({\rm Z}(\widetilde{Q}+\zeta,{\rm Ext}(\widetilde{S},\widetilde{R}{\langle}\zeta{\rangle}))+b_{0}({\rm Z}(\widetilde{Q}-\zeta,{\rm Ext}(\widetilde{S},\widetilde{R}{\langle}\zeta{\rangle}))$
    where $\zeta$ is a new infinitesimal (Lemma 4.8).
  2. (b)
    The real algebraic sets
    ${\rm Z}(\widetilde{Q}\pm\zeta,{\rm Ext}(\widetilde{S},\widetilde{R}{\langle}\zeta{\rangle})$
    if non-empty, are of dimension $p-1$, and satisfy the local dimension condition and are of degree $2Dd^{i}$. If $\pi$ denotes a generic projection to a real subspace of dimension $p$, then $\pi({\rm Z}(\widetilde{Q}\pm\zeta,{\rm Ext}(\widetilde{S},\widetilde{R}{\langle}\zeta{\rangle}))$ is a real hypersurface of degree $2Dd^{i}$ and is defined by a real polynomial of that degree. Denoting $\widetilde{Z}_{\pm\zeta}={\rm Z}(\widetilde{Q}\pm\zeta,{\rm Ext}(\widetilde{S},\widetilde{R}{\langle}\zeta{\rangle}))$ we prove that $b_{0}(\widetilde{Z})$ is bounded by $b_{0}(\pi(\widetilde{Z})-\mathrm{Sing}(\pi(\widetilde{Z})))$ (Lemma 4.10). Since, $\mathrm{Sing}(\pi(\widetilde{Z})))$ is a real algebraic subset of $\pi(\widetilde{Z})$, we can define its real points by taking the sum of squares of all partial derivatives of the polynomial defining $\pi(\widetilde{Z})$. We then use the Oleĭnik-Petrovskiĭ inequality [34] bounding the Betti numbers of real algebraic sets to bound the last quantity (Lemma 4.11).
Theorem 2 now follows after adding up the contributions from the different algebraic sets considered (namely, (a), (b) and (c) in the proof).
4.2.4. Proof of Theorem 2
Let
$\mathcal{P}=\{P_{1},\ldots,P_{s}\}\subset\mathrm{R}[X_{1},\ldots,X_{N}].$
We denote
$\widetilde{\mathcal{P}}=\bigcup_{i=1}^{s}\{P_{i}\pm{\varepsilon}\gamma_{i},P_{i}\pm\delta\gamma_{i}\}\subset\widetilde{\mathrm{R}}[X_{1},\ldots,X_{N}]\cup\{P_{0}\},$
where
$P_{0}=\gamma_{0}(\sum_{i=1}^{N}X_{i}^{2})-1,$
and
$\widetilde{\mathrm{R}}=\mathrm{R}{\langle}\gamma_{0},\gamma_{1},\ldots,\gamma_{s},{\varepsilon},\delta{\rangle}.$
We will use the following lemmas.
Lemma 4.4.
Let $\sigma\in\Sigma(\mathcal{P},V(\mathrm{R}))$ and $C\in\mathrm{Cc}({\mathcal{R}}(\sigma,V(\mathrm{R})))$. Then there exists a unique element $\widetilde{C}_{\sigma}\in\mathrm{Cc}({\mathcal{R}}(\widetilde{\phi}_{\sigma},{\rm Ext}(V(\mathrm{R}),\widetilde{\mathrm{R}}))),$ where
$\widetilde{\phi}_{\sigma}$ ::=:= $(P_{0}\leq 0)\wedge$
$\bigwedge_{i,\sigma(P_{i})=0}(-\delta\gamma_{i}\leq P_{i}\leq\delta\gamma_{i})\wedge$
$\bigwedge_{i,\sigma(P_{i})=1}(P_{i}-{\varepsilon}\gamma_{i}\geq 0)\wedge$
$\bigwedge_{i,\sigma(P_{i})=-1}(P+{\varepsilon}\gamma_{i}\leq 0),$
such that
$\widetilde{C}_{\sigma}\cap V(\mathrm{R})=C\cap V(\mathrm{R}).$
Proof.
Let $x\in C$. Then, $x\in{\mathcal{R}}(\widetilde{\phi}_{\sigma},{\rm Ext}(V(\mathrm{R}),\widetilde{\mathrm{R}}))$, and let $\widetilde{C}_{\sigma}$, be the semi-algebraically connected component of ${\mathcal{R}}(\widetilde{\phi}_{\sigma},{\rm Ext}(V(\mathrm{R}),\widetilde{\mathrm{R}}))$ containing $x$. Then $\widetilde{C}_{\sigma}\cap V(\mathrm{R})=C\cap V(\mathrm{R})$ (see [7, Proposition 13.7]). ∎
Lemma 4.5.
For $I\subset\{1,\ldots,s\}$ and $\alpha:I\rightarrow\{\pm{\varepsilon},\pm\delta\}$,
$\dim\left(V\cap\bigcap_{i\in I}{\rm Z}(P_{i}+\alpha(i)\gamma_{i},\mathrm{C}^{N})\right)=p-\mathrm{card}(I).$
Proof.
Follows from the fact that $\alpha(i)\gamma_{i},i\in I$ are independent transcendentals over $\mathrm{C}$. ∎
Remark 25.
Note that as an immediate consequence of Lemma 4.5 we get that
$V\cap\bigcap_{i\in I}{\rm Z}(P_{i}+\alpha(i)\gamma_{i},\mathrm{C}^{N})=\emptyset,$
whenever $\mathrm{card}(I)>p$.
Lemma 4.6.
Let $V\subset\mathrm{R}^{N}$ be a closed semi-algebraic set and $\mathcal{Q}\subset\mathrm{R}[X_{1},\ldots,X_{N}]$ a finite subset. Let $C\in\mathrm{Cc}(S)$, where $S\subset V$ is the semi-algebraic set defined by
$S=\{x\in V\mid\bigwedge_{Q\in\mathcal{Q}}(Q(x)\geq 0)\}.$
Then there exists a subset $\mathcal{Q}^{\prime}\subset\mathcal{Q}$, and $D\in\mathrm{Cc}(V\cap{\rm Z}(\mathcal{Q}^{\prime},\mathrm{R}^{N}))$ such that $D\subset C$.
Proof.
Let $\mathcal{Q}^{\prime}$ be a maximal subset of $\mathcal{Q}$ with the property that ${\rm Z}(\mathcal{Q}^{\prime},\mathrm{R}^{N})\cap C\neq\emptyset$. Then $\mathcal{Q}^{\prime}$ satisfies the required property in the lemma (see [7, Proposition 13.1]). ∎
Lemma 4.7.
Let $\sigma\in\{0,1,-1\}^{\mathcal{P}}$ and $C\in\mathrm{Cc}({\mathcal{R}}(\sigma,V(\mathrm{R})))$. Then there exist $I\subset\{1,\ldots,s\}$, $\alpha:I\rightarrow\{\pm{\varepsilon},\pm\delta\}$, and $D$ in the union of
$\mathrm{Cc}\left(\bigcap_{i\in I}{\rm Z}(P_{i}+\alpha(i)\gamma_{i},{\rm Ext}(V(\mathrm{R}),\widetilde{\mathrm{R}}))\right)$
and
$\mathrm{Cc}\left({\rm Z}(P_{0}-{\varepsilon}\gamma_{0},{\rm Ext}(V(\mathrm{R}),\widetilde{\mathrm{R}}))\cap\bigcap_{i\in I}{\rm Z}(P_{i}+\alpha(i)\gamma_{i},{\rm Ext}(V(\mathrm{R}),\widetilde{\mathrm{R}}))\right),$
such that
$\alpha(i)=\begin{cases}{\varepsilon}\text{ if }\sigma(P_{i})=1,\\ -{\varepsilon}\text{ if }\sigma(P_{i})=-1,\\ \delta\text{ or }-\delta\text{ if }\sigma(P_{i})=0,\end{cases}$
and
$D\subset\widetilde{C}_{\sigma}.$
Proof.
Follows from Lemmas 4.4 and 4.6. ∎
Lemma 4.8.
Let $S\subset\mathrm{R}^{N}$ be a closed and bounded semi-algebraic set and $Q\in\mathrm{R}[X_{1},\ldots,X_{N}]$, such that at every point $x\in{\rm Z}(Q,S)$, $\dim_{x}{\rm Z}(Q,S)<\dim_{x}S$. Then,
$b_{0}({\rm Z}(Q,S))\leq b_{0}({\rm Z}(Q-\zeta,{\rm Ext}(S,\mathrm{R}{\langle}\zeta{\rangle})))+b_{0}({\rm Z}(Q+\zeta,{\rm Ext}(S,\mathrm{R}{\langle}\zeta{\rangle}))).$
Proof.
Let $\widetilde{S}$ be the intersection of ${\rm Ext}(S,\mathrm{R}{\langle}\zeta{\rangle})$ with the set defined by $-\zeta\leq Q\leq\zeta$. Let $C\in\mathrm{Cc}({\rm Z}(Q,S))$ and $x\in C$. Then, it follows from the hypothesis that $\dim_{x}{\rm Z}(Q,C)<\dim_{x}S$. Hence, for all $r>0$, denoting (here and elsewhere) by $B_{N}(x,r)$ the open Euclidean ball centered at $x$ of radius $r$, ${\rm Z}(Q,C)\cap B_{N}(x,r)$ is non-empty and properly contained in $S\cap B_{N}(x,r)$. So $S-{\rm Z}(Q,C)\neq\emptyset$ and $x\in\overline{S-{\rm Z}(Q,C)}$. Applying the semi-algebraic curve selection lemma (see for example [7, Theorem 3.19]), there exists a semi-algebraic curve $\gamma:[0,1]\rightarrow S$, such that $\gamma(0)=x$, and $\gamma((0,1])\subset S-{\rm Z}(Q,C)$. In particular, $Q(\gamma(1))\neq 0$.
Suppose without loss of generality that $Q(\gamma(1))>0$. Using the intermediate value property of the field $\mathrm{R}{\langle}\zeta{\rangle}$ ([7, Proposition 4.3]), there exists $\widetilde{t}\in\mathrm{R}{\langle}\zeta{\rangle},\widetilde{t}>0$, such that $Q({\rm Ext}(\gamma,\mathrm{R}{\langle}\zeta{\rangle})(\widetilde{t}))=\zeta$, and let $\widetilde{x}={\rm Ext}(\gamma,\mathrm{R}{\langle}\zeta{\rangle})(\widetilde{t})$. Then $\widetilde{x}\in{\rm Z}(Q-\zeta,{\rm Ext}(S,\mathrm{R}{\langle}\zeta{\rangle}))$. Let $\widetilde{C}\in\mathrm{Cc}({\rm Z}(Q-\zeta,{\rm Ext}(S,\mathrm{R}{\langle}\zeta{\rangle})))$ with $\widetilde{x}\in\widetilde{C}$. Then, $\lim_{\zeta}\widetilde{C}\subset C$. Thus we have proved that for each $C\in\mathrm{Cc}({\rm Z}(Q,S))$, there exists $\widetilde{C}\in\mathrm{Cc}({\rm Z}(Q-\zeta,{\rm Ext}(S,\mathrm{R}{\langle}\zeta{\rangle})))\cup\mathrm{Cc}({\rm Z}(Q+\zeta,{\rm Ext}(S,\mathrm{R}{\langle}\zeta{\rangle})))$, with $\lim_{\zeta}\widetilde{C}\subset C$. The lemma follows. ∎
Lemma 4.9.
Let $V\subset\mathrm{C}^{N}$ be a variety with $\dim V=p$. Then for all generically chosen linear maps $\pi:\mathrm{C}^{N}\rightarrow\mathrm{C}^{p+1}$, and set of points $w\in W=\pi(V)$, such that $\mathrm{card}(\pi^{-1}(w)\cap V)=1$ is Zariski dense in $W$.
Proof.
It follows from the Noether normalization theorem (see for example [15, Theorem 2.2.9]) that for a generic linear projection map $P:\mathrm{C}^{N}\rightarrow\mathrm{C}^{p}$ of rank $p$, $V_{\ell}:=P^{-1}(\ell)\cap V$ is (non-empty and) finite for all $\ell\in L=\mathrm{Im}(P)\subset\mathrm{C}^{N}$.
Let $P_{0}:\mathrm{C}^{N}\rightarrow\mathrm{C}^{N}$ be a generic projection of rank $p$, and denote $L_{0}=\mathrm{Im}(P_{0})\subset\mathrm{C}^{N}$.
The set $\mathcal{Q}$ of linear projections $Q:\mathrm{C}^{N}\rightarrow\mathrm{C}^{N}$ of rank $p+1$ such that $L_{0}\subset\mathrm{Im}(Q)$ is a subvariety of the variety of all projections of rank $p+1$.
Let
$\Gamma=\{\ell,Q)\in L_{0}\times\mathcal{Q}\mid Q|_{V_{\ell}}\text{is not injective}\}.$
Notice that $\dim\Gamma<p+\dim\mathcal{Q}$.
Let $\pi_{1}:\Gamma\rightarrow U$, $\pi_{2}:\Gamma\rightarrow\mathcal{Q}$ be the two projection maps.
For a generic choice of $Q_{0}\in\mathcal{Q}$,
(4.2) $\dim\pi_{2}^{-1}(Q_{0})<p,$
since otherwise $\dim\Gamma=p+\dim\mathcal{Q}$.
Let $W^{\prime}=Q_{0}(P_{0}^{-1}(\pi_{1}(\pi_{2}^{-1}(Q_{0}))))$. Then $\dim W^{\prime}<p$ (using (4.2) and the fact that $P_{0}|_{V}$ is a finite map).
Now define $\pi:\mathrm{C}^{N}\rightarrow\mathrm{Im}(Q_{0})\cong\mathrm{C}^{p+1}$, by $\pi(x)=Q_{0}(x)$. Then, with $W=\pi(V)$, $W^{\prime}\subset W$, with $\dim(W^{\prime})<p=\dim(W)$. Hence, we have for each $w\in W-W^{\prime}$, $\mathrm{card}(\pi^{-1}(w)\cap V)=1$. This proves the lemma. ∎
Lemma 4.10.
Let $V\subset\mathrm{C}^{N}$ be a variety with $\dim V=p$. Then for all generically chosen linear maps $\pi:\mathrm{C}^{N}\rightarrow\mathrm{C}^{p+1}$, $W=\pi(V)\subset\mathrm{C}^{p+1}$ is a hypersurface, and the subset $W^{\prime}\subset W$, defined by $W^{\prime}=\{w\in W\mid\mathrm{card}(\pi^{-1}(w)\cap V)>1\}$ is contained in $\mathrm{Sing}(W)$.
Proof.
It follows from Noether normalization theorem (see for example [15, Theorem 2.2.9]) that for $\pi$ generic, $\pi(V)$ is a hypersurface in $\mathrm{C}^{p+1}$.
Let $w\in W=\pi(V)$, and let $V_{w}=\pi^{-1}(w)\cap V$. It follows from Lemma 4.9, that there exists a neighborhood $U_{w}\subset W$ of $w$ in $W$, such that for each $z\in V_{w}$ has a neighborhood $U_{z}$ in $\pi^{-1}(U_{w})\cap V$, such that $\pi|_{U_{z}}$ is injective, and moreover for $z\neq z^{\prime}\in V_{w}$ $U_{z}$ and $U_{z^{\prime}}$ are disjoint.
It now follows that
$\mathrm{mult}_{w}(W)\geq\mathrm{card}(V_{w}).$
This implies that if $\mathrm{card}(V_{w})>1$, then $\mathrm{mult}_{w}(W)>1$, and it follows from Lemma 4.3 that in this case $w\in\mathrm{Sing}(W)$.
Lemma 4.11.
Let $V\subset\mathrm{C}^{N}$ be a real variety such that $\dim_{x}V(\mathrm{R})=p$ for all $x\in V(\mathrm{R})$. Then, for all generically chosen linear maps $\pi:\mathrm{R}^{N}\rightarrow\mathrm{R}^{p+1}$,
$b_{0}(V(\mathrm{R}))\leq b_{0}(\pi_{\mathrm{C}}(V)(\mathrm{R})-\mathrm{Sing}(\pi_{\mathrm{C}}(V))),$
where $\pi_{\mathrm{C}}:\mathrm{C}^{N}\rightarrow\mathrm{C}^{p+1}$ obtained from $\pi$ by tensoring with $\mathrm{C}$.
Proof.
It follows from Lemma 4.10, that for any $z\in\pi_{\mathrm{C}}(V)-\mathrm{Sing}(\pi_{\mathrm{C}}(V))$, $\mathrm{card}(\pi_{\mathrm{C}}^{-1}(z)\cap V)=1$.
Now, since $V$ is defined over $\mathrm{R}$, for $x\in\pi(V(\mathrm{R}))$, $\mathrm{card}(\pi_{\mathrm{C}}^{-1}(x)\cap V)=1$ implies that $\pi_{\mathrm{C}}^{-1}(x)\cap V\subset V(\mathrm{R})$, since otherwise $\pi_{\mathrm{C}}^{-1}(x)\cap V$ will contain at least one pair of conjugate points $z\neq\bar{z},z,\bar{z}\in\pi_{\mathrm{C}}^{-1}(x)\cap V$.
Taken together the previous two statements imply:
$\pi_{\mathrm{C}}^{-1}(\pi_{\mathrm{C}}(V)(\mathrm{R})-\mathrm{Sing}(\pi_{\mathrm{C}}(V)))\cap V\subset V(\mathrm{R}).$
We now prove that for each $C\in\mathrm{Cc}(V(\mathrm{R}))$, there exists $D\in\mathrm{Cc}(\pi_{\mathrm{C}}(V)(\mathrm{R})-\mathrm{Sing}(\pi_{\mathrm{C}}(V)))$, such that $\pi^{-1}(D)\subset C$. It would follow that
$b_{0}(V(\mathrm{R}))\leq b_{0}(\pi_{\mathrm{C}}(V)(\mathrm{R})-\mathrm{Sing}(\pi_{\mathrm{C}}(V))).$
Let $C\in\mathrm{Cc}(V(\mathrm{R}))$. Since $\pi(V(\mathrm{R}))\subset\pi_{\mathrm{C}}(V)(\mathrm{R})$there exists $D^{\prime}\in\mathrm{Cc}(\pi_{\mathrm{C}}(V)(\mathrm{R}))$ such that $\pi(C)\subset D^{\prime}$.
Since the local real dimension of $V(\mathrm{R})$ is equal to $p$ at all points of $V(\mathrm{R})$, $\dim_{\mathrm{R}}C=\dim_{\mathrm{R}}D^{\prime}=p$. Moreover, $\dim_{\mathrm{R}}D^{\prime}\cap\mathrm{Sing}(\pi_{C}(V))<p$, and hence $D^{\prime}-\mathrm{Sing}(\pi_{C}(V))\neq\emptyset$. Now let $D\in\mathrm{Cc}(D^{\prime}-\mathrm{Sing}(\pi_{C}(V)))\subset\mathrm{Cc}(\pi_{\mathrm{C}}(V)(\mathrm{R})-\mathrm{Sing}(\pi_{\mathrm{C}}(V)))$. Then, $\pi^{-1}(D)\subset C$ and we are done. ∎
Lemma 4.12.
Let $V={\rm Z}(P,\mathrm{C}^{m})$ and $P\in\mathrm{R}[X_{1},\ldots,X_{m}]_{\leq d}$. Then,
$b_{0}(V(\mathrm{R})-\mathrm{Sing}(V))\leq 14\cdot d\cdot(4d-1)^{m-1}.$
Proof.
Let $Q=\sum_{i=1}^{m}\left(\frac{\partial P}{\partial X_{i}}\right)^{2}$. Then, there exists $R>0$, such that
$b_{0}(V(\mathrm{R})-\mathrm{Sing}(V))=b_{0}((V(\mathrm{R})-\mathrm{Sing}(V))\cap B_{m}(0,R)),$
and
$b_{0}((V(\mathrm{R})-\mathrm{Sing}(V))\cap B_{m}(0,R))=b_{0}(S)$
where $S\subset\mathrm{R}{\langle}{\varepsilon}{\rangle}^{N}$ is defined by
$(P=0)\wedge(Q\geq{\varepsilon})\wedge(\sum_{i=1}^{m}X_{i}^{2}-R\leq 0).$
It follows from Lemma 4.6 that $b_{0}(S)$ is bounded by the number of semi-algebraically connected components of the $\binom{3}{1}+\binom{3}{2}+\binom{3}{3}=7$ algebraic sets defined by taking one, two or three polynomials $P,Q-{\varepsilon},\sum_{i=1}^{m}X_{i}^{2}-R$ at a time.
The number of semi-algebraically connected components of each one them is bounded by
$2d(4d-1)^{m-1},$
using the basic inequality on the Betti numners of real algebraic sets due to Oleĭnik and Petrovskiĭ [34] (see for example [7, Proposition 7.28]). The lemma follows. ∎
Proof of Theorem 2.
First of all observe that using the fact that
$\deg(V)=\sum_{i=1}^{m}\deg(V_{i}),$
and for each $\sigma\in\{0,1,-1\}^{\mathcal{P}}$,
$b_{0}({\mathcal{R}}(\sigma,V(\mathrm{R})))\leq\sum_{i=1}^{m}b_{0}({\mathcal{R}}(\sigma,V_{i}(\mathrm{R}))),$
we can reduce to the case $m=1$ and assume that $V=V_{1}$.
Using Lemma 4.7 we are reduced to bounding the number of semi-algebraically connected components of the real algebraic sets,
$\bigcap_{i\in I}{\rm Z}(P_{i}+\alpha(i)\gamma_{i},{\rm Ext}(V(\mathrm{R}),\widetilde{\mathrm{R}})),$
and
${\rm Z}(P_{0}-{\varepsilon}\gamma_{0},{\rm Ext}(V(\mathrm{R}),\widetilde{\mathrm{R}}))\cap\bigcap_{i\in I}{\rm Z}(P_{i}+\alpha(i)\gamma_{i},{\rm Ext}(V(\mathrm{R}),\widetilde{\mathrm{R}})),$
where $I\subset[1,s]$, and $\alpha$ as defined in Lemma 4.7. Using Lemma 4.5 and Remark 25, we only need to consider $I$, with $\mathrm{card}(I)\leq p$ for the algebraic sets $\bigcap_{i\in I}{\rm Z}(P_{i}+\alpha(i)\gamma_{i},{\rm Ext}(V(\mathrm{R}),\widetilde{\mathrm{R}}))$ and of cardinality at most $p-1$ for the algebraic sets ${\rm Z}(P_{0}-{\varepsilon}\gamma_{0},{\rm Ext}(V(\mathrm{R}),\widetilde{\mathrm{R}}))\cap\bigcap_{i\in I}{\rm Z}(P_{i}+\alpha(i)\gamma_{i},{\rm Ext}(V(\mathrm{R}),\widetilde{\mathrm{R}}))$.
Thus the total number of algebraic sets we need to consider is bounded by
$\sum_{0\leq j\leq p}4^{j}\binom{s}{j}+\sum_{0\leq j\leq p-1}4^{j}\binom{s}{j}=2\cdot\sum_{0\leq j\leq p-1}4^{j}\binom{s}{j}+4^{p}\binom{s}{p}.$
Each one of these algebraic sets is the intersection of ${\rm Ext}(V(\mathrm{R}),\widetilde{\mathrm{R}})$ with an real algebraic set defined by a polynomial of degree at most $2d$ of the form
$Q_{I,\alpha}=\sum_{i\in I}(P_{i}+\alpha(i)\gamma_{i})^{2},$
or
$Q_{I,\alpha,0}=(P_{0}-{\varepsilon}\gamma_{0})^{2}+\sum_{i\in I}(P_{i}+\alpha(i)\gamma_{i})^{2}.$
Using Lemma 4.8 it suffices to bound the number of semi-algebraically connected components of ${\rm Z}(Q_{I,\alpha}\pm\zeta,{\rm Ext}(V(\mathrm{R}),\widetilde{\mathrm{R}}{\langle}\zeta{\rangle}))$ and ${\rm Z}(Q_{I,\alpha,0}\pm\zeta,{\rm Ext}(V(\mathrm{R}),\widetilde{\mathrm{R}}{\langle}\zeta{\rangle}))$.
We accomplish this using Lemmas 4.11 and 4.12.
Note that for any $I\subset[1,s]$ ${\rm Z}(Q_{I,\alpha,0}\pm\zeta,{\rm Ext}(V,\widetilde{\mathrm{R}}{\langle}\zeta{\rangle}))$ and ${\rm Z}(Q_{I,\alpha}\pm\zeta,{\rm Ext}(V,\widetilde{\mathrm{R}}{\langle}\zeta{\rangle}))$ are real algebraic sets in $\widetilde{R}{\langle}\zeta{\rangle}[i]^{N}$ of dimension $p-1$ if non-empty, satisfying the property of constancy of local real dimension, and degree bounded by $2Dd$. If $I$ is empty, $V$ is of dimension $p$ and degree $D$ and satisfies the property of constancy of local real dimension.
Thus in order to bound $\sum_{\sigma\in\{0,1,-1\}^{\mathcal{P}}}b_{0}({\mathcal{R}}(\sigma,V(\mathrm{R})))$ it suffices to bound
  1. (a)
    $b_{0}({\rm Z}(Q_{I,\alpha,0}\pm\zeta,{\rm Ext}(V,\widetilde{\mathrm{R}}{\langle}\zeta{\rangle})))$, $I\subset[1,s],1\leq\mathrm{card}(I)<p$,
  2. (b)
    $b_{0}({\rm Z}(Q_{I,\alpha}\pm\zeta,{\rm Ext}(V,\widetilde{\mathrm{R}}{\langle}\zeta{\rangle})))$, $I\subset[1,s],1\leq\mathrm{card}(I)\leq p$, and
  3. (c)
    $b_{0}(V(\mathrm{R}))$.
We apply Lemmas 4.11 and 4.12 to each of the three cases listed above, and add their contributions to obtain
$\sum_{\sigma\in\{0,1,-1\}^{\mathcal{P}}}b_{0}({\mathcal{R}}(\sigma,V(\mathrm{R})))$$\leq$ $\sum_{j=1}^{p}4^{j}\binom{s}{j}14\cdot 2Dd(8Dd-1)^{p-1}+$
$\sum_{j=1}^{p-1}4^{j}\binom{s}{j}14\cdot 2Dd(8Dd-1)^{p-1}+$
$14\cdot D(4D-1)^{p}$
$\leq$ $O(D)^{p}\cdot\left(\sum_{j=1}^{p}4^{j}\binom{s}{j}d^{p}+D\right).$
We now prove Theorem 3 by modifying the proof of Theorem 2 in the special case when $\mathcal{P}=\emptyset$. The constancy of the local dimension of $V$ is used in the proof of Theorem 2 only in Lemma 4.11. The following lemma replaces Lemma 4.11 in the proof of Theorem 3.
Lemma 4.13.
Let $V\subset\mathrm{C}^{N}$ be an algebraic subset of dimension $p$ defined over $\mathrm{R}$. Then, for all generically chosen linear maps $\pi:\mathrm{R}^{N}\rightarrow\mathrm{R}^{p+1}$,
$b_{0}(V(\mathrm{R})_{p})\leq b_{0}(\pi_{\mathrm{C}}(V)(\mathrm{R})-\mathrm{Sing}(\pi_{\mathrm{C}}(V))),$
where $\pi_{\mathrm{C}}:\mathrm{C}^{N}\rightarrow\mathrm{C}^{p+1}$ obtained from $\pi$ by tensoring with $\mathrm{C}$.
Proof.
We follow closely the proof of Lemma 4.11. Using the same argument as in the proof of Lemma 4.11, we have
$\pi_{\mathrm{C}}^{-1}(\pi_{\mathrm{C}}(V)(\mathrm{R})-\mathrm{Sing}(\pi_{\mathrm{C}}(V)))\cap V\subset V(\mathrm{R}).$
We now prove that for each $C\in\mathrm{Cc}(V(\mathrm{R})_{p})$, there exists $D\in\mathrm{Cc}(\pi_{\mathrm{C}}(V)(\mathrm{R})-\mathrm{Sing}(\pi_{\mathrm{C}}(V)))$, such that $\pi^{-1}(D)\subset C$. It would follow that
$b_{0}(V(\mathrm{R})_{p})\leq b_{0}(\pi_{\mathrm{C}}(V)(\mathrm{R})-\mathrm{Sing}(\pi_{\mathrm{C}}(V))).$
Let $C\in\mathrm{Cc}(V(\mathrm{R})_{p})$. Since $\pi(V(\mathrm{R}))\subset\pi_{\mathrm{C}}(V)(\mathrm{R})$ there exists $D^{\prime}\in\mathrm{Cc}(\pi_{\mathrm{C}}(V)(\mathrm{R}))$ such that $\pi(C)\subset D^{\prime}$.
Since the local real dimension of $V(\mathrm{R})_{p}$ is equal to $p$ at all points of $V(\mathrm{R})_{p}$, $\dim_{\mathrm{R}}C=\dim_{\mathrm{R}}D^{\prime}=p$. Moreover, since $\dim_{\mathrm{R}}(D^{\prime}\cap\mathrm{Sing}(\pi_{C}(V)))<p$, $D^{\prime}-\mathrm{Sing}(\pi_{C}(V))\neq\emptyset$. Now let $D\in\mathrm{Cc}(D^{\prime}-\mathrm{Sing}(\pi_{C}(V)))\subset\mathrm{Cc}(\pi_{\mathrm{C}}(V)(\mathrm{R})-\mathrm{Sing}(\pi_{\mathrm{C}}(V)))$. Then, $\pi^{-1}(D)\subset C$ and we are done. ∎
Proof of Theorem 3.
The theorem follows from Lemma 4.13 and Lemma 4.12. ∎
Proof of Theorem 4.
Observe that if $H\subset\mathbb{P}_{C}^{N}$ is a generically chosen real hyperplane, then
$\sum_{\sigma\in\{0,1,-1\}^{\mathcal{P}}}b_{0}({\mathcal{R}}(\sigma,V(\mathrm{R})))\leq\sum_{\sigma\in\{0,1,-1\}^{\mathcal{P}}}b_{0}({\mathcal{R}}(\pi,V(\mathrm{R})\setminus H))+\sum_{\sigma\in\{0,1-1\}^{\mathcal{P}}}b_{0}({\mathcal{R}}(\pi,(V\cap H)(\mathrm{R}))).$
We can bound the first term on the right using the affine bound (i.e. Theorem 2) and proceed by induction to bound the second term noting that $\dim V\cap H=p-1$ and $(V\cap H)(\mathrm{R})$ satisfies the property of constancy of local real dimension. ∎
4.3. Proofs of Theorems 5 and 6
We can assume without loss of generality that $V$ is a complete intersection variety having degree $D$ and dimension $p$. If the $V$ is a union of complete intersection varieties, $V_{1},\ldots,V_{m}$, then the sum of the degrees of the of $V_{i}$ is bounded by $D$, and
$\sum_{\sigma\in\{0,1,-1\}^{\mathcal{P}}}b_{0}({\mathcal{R}}(\sigma,V(\mathrm{R})))\leq\sum_{i=1}^{m}\sum_{\sigma\in\{0,1,-1\}^{\mathcal{P}}}b_{0}({\mathcal{R}}(\sigma,V_{i}(\mathrm{R}))).$
Thus, it would suffice to prove the inequality in the proposition for each $V_{i}$ and then take the sum of the upper bounds.
Henceforth $V$ is a complete intersection variety in $\mathrm{C}^{N}$ of degree $D$ and dimension $p$. Thus there exists polynomials, $Q_{1},\ldots,Q_{N-p}\in\mathrm{C}[Z_{1},\ldots,Z_{N}]$ such that $V$ is the set of common zeros in $\mathrm{C}^{N}$ of $Q_{1},\ldots,Q_{N-p}$, and $D=d_{1}\cdots d_{n-p}$, where $d_{i}=\deg(Q_{i}),1\leq i\leq N-p$. We can also assume without loss of generality that $d_{1}\geq\cdots\geq d_{N-p}$. Let $q,1\leq q\leq N-p$ the largest integer such that $d_{q}>1$. Then $d_{1},\ldots,d_{q}\geq 2$ and $d_{q+1}=\cdots=d_{n-p}=1$. So $D=d_{1}\cdots d_{q}\geq 2^{q}$. Hence, $q\leq\log_{2}D$. Thus, $V$ is contained in an affine subspace in $\mathrm{C}^{N}$ of dimension $N-(N-p-q)=p+q\leq p+\log_{2}D$. We can pass to this subspace if needed and assume that $N\leq p+\log_{2}D$.
4.3.1. Outline of the proof of Theorem 5
There are several new ingredients. Firstly, as in the proof of Theorem 2, adding a new inequality (that defining a ball of large radius) we can restrict our attention to the realizations of sign conditions which are bounded. Then using infinitesimal deformations of the kind already used in the proof of Theorem 2, we go to a situation where the perturbed polynomials are in general position (meaning that their zeros in $\mathbb{P}^{N}_{\mathrm{C}}$ are non-singular complete intersections), and for each semi-algebraically connected component $C$ of the realization on $V(\mathrm{R})$ of any realizable sign condition on the original family, there exists a semi-algebraically connected component $\widetilde{C}$ of the realization on the extension of $V(\mathrm{R})$ of a strict sign condition on the perturbed family such that $\widetilde{C}$ is semi-algebraically homotopy equivalent to the extension of $C$ (Lemma 4.14). This is stronger than just requiring $\widetilde{C}\cap\mathrm{R}^{N}=C$ (as in the proof of Theorem 2), but we need the stronger property in this proof.
We then observe that each bounded semi-algebraically connected component $C$ of a realizable sign condition of $\mathcal{P}$ on $V(\mathrm{R})$ is semi-algebraically homotopically equivalent to a semi-algebraically connected component of the complement in $\overline{V}(\mathrm{R})$ of the union of the set of zeros of the homogenizations of the perturbed family (Lemma 4.15).
Now using the fact that $\overline{V}(\mathrm{R})$ is non-singular (and hence a compact semi-algebraic manifold), and Lefschetz duality, in order to bound the sum of all the Betti numbers of the realizations on $V(\mathrm{R})$ of all realizable sign conditions on the original family, it suffices to bound the sum of the Betti numbers of $V$ and the those of the union of the algebraic sets defined by the perturbed family (Lemma 4.16).
At this point we use a standard reduction via the Mayer-Vietoris spectral sequence to reduce this latter problem to bounding the Betti numbers of the algebraic sets on $\overline{V}(\mathrm{R})$ defined by the perturbed family (Lemma 4.17).
These intersections are non-singular complete intersections. For the corresponding algebraic sets over $\mathrm{C}$ whch are projective non-singular complete intersections, one can bound the Betti numbers using a classical formula coming from Chern class computation of the tangent bundle (Lemma 4.19). This gives us an upper bound on the Betti numbers of the real parts of these varieties using the Thom-Smith inequalities (Lemma 4.18). The theorem follows by aggregrating the contributions to the Betti numbers from the different algebraic sets.
4.3.2. Proof of Theorem 5
We will need the following notation and lemmas.
Notation 4.5.
For any closed and bounded semi-algebraic set $S$ we denote by $b(S)=\sum_{i\geq 0}b_{i}(S)$ the sum of the $\mathbb{Z}/2\mathbb{Z}$-Betti numbers of $S$.
Notation 4.6.
Let $\mathcal{P}=\{P_{1},\ldots,P_{s}\}\subset\mathrm{R}[X_{1},\ldots,X_{N}]$ be a finite set. Let $H_{1},\ldots,H_{s}\in\mathrm{R}[X_{1},\ldots,X_{N}]$ with $H_{1}(x),\ldots,H_{N}(x)>0$, for all $x\in S$, and
$\widetilde{\mathcal{P}}=\bigcup_{i=1}^{s}\{P_{i}\pm{\varepsilon}H_{i},P_{i}\pm\delta H_{i}\}\subset\mathrm{R}{\langle}{\varepsilon},\delta{\rangle}[X_{1},\ldots,X_{N}].$
For $\sigma\in\sigma\in\{0,1,-1\}^{\mathcal{P}}$ define $\widetilde{\sigma}\in\{0,1,-1\}^{\widetilde{\mathcal{P}}}$ as follows. For each $i,1\leq i\leq s$, If $\sigma(P_{i})=0$:
$\widetilde{\sigma}(P_{i}+{\varepsilon}H_{i})$ $=$ $+1,$
$\widetilde{\sigma}(P_{i}-{\varepsilon}H_{i})$ $=$ $-1,$
$\widetilde{\sigma}(P_{i}+\delta H_{i})$ $=$ $+1,$
$\widetilde{\sigma}(P_{i}-\delta\ H_{i})$ $=$ $-1.$
If $\sigma(P_{i})=1$:
$\widetilde{\sigma}(P_{i}+{\varepsilon}H_{i})$ $=$ $1,$
$\widetilde{\sigma}(P_{i}-{\varepsilon}H_{i})$ $=$ $1,$
$\widetilde{\sigma}(P_{i}+\delta H_{i})$ $=$ $1,$
$\widetilde{\sigma}(P_{i}-\delta\ H_{i})$ $=$ $1$
If $\sigma(P_{i})=-1$:
$\widetilde{\sigma}(P_{i}+{\varepsilon}H_{i})$ $=$ $-1,$
$\widetilde{\sigma}(P_{i}-{\varepsilon}H_{i})$ $=$ $-1,$
$\widetilde{\sigma}(P_{i}+\delta H_{i})$ $=$ $-1,$
$\widetilde{\sigma}(P_{i}-\delta\ H_{i})$ $=$ $-1.$
Lemma 4.14.
Let $S\subset\mathrm{R}^{N}$ be a closed and bounded semi-algebraic set. Then using the notation in Notation 4.6, for each $i\geq 0$,
$\sum_{\sigma\in\{0,1,-1\}^{\mathcal{P}}}b_{i}({\mathcal{R}}(\sigma,S))\leq\sum_{\widetilde{\sigma}\in\{1,-1\}^{\widetilde{\mathcal{P}}}}b_{i}({\mathcal{R}}(\widetilde{\sigma},{\rm Ext}(S,\mathrm{R}{\langle}{\varepsilon},\delta{\rangle}))).$
Proof.
Using the fact that $S$ is closed and bounded it is easy to show that ${\rm Ext}({\mathcal{R}}(\sigma,S),\mathrm{R}{\langle}{\varepsilon},\delta{\rangle}$ is semi-algebraically homotopy equivalent to ${\mathcal{R}}(\widetilde{\sigma},{\rm Ext}(S,\mathrm{R}{\langle}{\varepsilon},\delta{\rangle}))$. ∎
Lemma 4.15.
Let $\overline{V}\subset\mathbb{P}_{\mathrm{C}}^{N}$ and $V(\mathrm{R})=\overline{V}\cap\mathrm{R}^{N}$, where $\overline{V}$ variety. Let $\mathcal{P}=\{P_{1},\ldots,P_{s}\}\subset\mathrm{R}[X_{1},\ldots,X_{N}]$ be a finite subset, and let $P_{0}=\sum_{i=1}^{N}X_{i}^{2}-R$ with $R\in\mathrm{R},R>0$. Let for $i=0,\ldots,s$, $H_{i}\in\mathrm{R}[X_{1},\ldots,X_{N}]$ be generic polynomials with $H_{i}(x)>0$ for all $x\in\mathrm{R}^{N}$, and $\deg(H_{i})=\deg(P_{i})$. Let $\widetilde{\mathcal{P}}=\bigcup_{i=0}^{s}\{P_{i}\pm{\varepsilon}H_{i},P_{i}\pm\delta H_{i}\}\subset\mathrm{R}{\langle}{\varepsilon},\delta{\rangle}[X_{1},\ldots,X_{N}]$.
Then, for all large $R>0$, and $i\geq 0$,
$\sum_{\sigma\in\{0,1,-1\}^{\mathcal{P}}}b_{i}({\mathcal{R}}(\sigma,V(\mathrm{R})))\leq\sum_{\widetilde{\sigma}\in\{1,-1\}^{\widetilde{\mathcal{P}}}}b_{i}({\mathcal{R}}(\widetilde{\sigma},{\rm Ext}(V(\mathrm{R}),\mathrm{R}{\langle}{\varepsilon},\delta{\rangle}))).$
Proof.
It follows from the conic structure at infinity of semi-algebraic sets [13] that for all large enough $R>0$, for each $\sigma\in\{0,1,-1\}^{\mathcal{P}}$, ${\mathcal{R}}(\sigma,V(\mathrm{R}))$ is semi-algebraically homeomorphic to ${\mathcal{R}}(\sigma,V(\mathrm{R})\cap B_{N}(0,R))$. Let $\sigma_{0}$ denote the extension of the sign condition $\sigma$ to $\mathcal{P}_{0}=\{P_{0}\}\cup\mathcal{P}$, by setting $\sigma_{0}(P_{0})=-1$. Then, for all large enough $R>0$, ${\mathcal{R}}(\sigma,V(R))$ is semi-algebraically homeomorphic to ${\mathcal{R}}(\sigma_{0},V(\mathrm{R}))$. Now let $S=V(\mathrm{R})\cap\overline{B_{N}(0,2R)}$. Notice that $S$ is a closed and bounded semi-algebraic set, and ${\mathcal{R}}(\sigma_{0},V(\mathrm{R}))={\mathcal{R}}(\sigma_{0},S)$. The lemma now follows from Lemma 4.14. ∎
In the following lemma we will consider a projective variety in $\mathbb{P}^{N}_{\mathrm{C}}$ with homogeneous coordinates $[X_{0}:\cdots:X_{N}]$. We will identify $\mathrm{C}^{N}$ (resp. $\mathrm{R}^{N}$) with the affine chart in $\mathbb{P}^{N}_{\mathrm{C}}$ (resp. $\mathbb{P}^{N}_{\mathrm{R}}$) defined by $X_{0}\neq 0$. For any polynomial $Q\in\mathrm{C}[X_{1},\ldots,X_{N}]$, we will denote by $Q^{h}\in\mathrm{C}[X_{0},\ldots,X_{N}]$ its homogenization.
Finally, for any semi-algebraic set $S\subset\mathrm{R}^{N}$, we denote by $\mathrm{Cc}_{b}(S)$, the subset of $\mathrm{Cc}(S)$, consisting of the semi-algebraically connected components of $S$ which are bounded.
Lemma 4.16.
Let $\overline{V}\subset\mathbb{P}_{\mathrm{C}}^{N}$ and $V(\mathrm{R})=\overline{V}\cap\mathrm{R}^{N}$, where $\overline{V}$ is a non-singular real variety. Let $\mathcal{Q}\subset\mathrm{R}[X_{1},\ldots,X_{N}]$ be a finite subset.
Then, for each $i,0\leq i\leq\dim\overline{V}$,
$\sum_{\sigma\in\{1,-1\}^{\mathcal{Q}}}\sum_{C\in\mathrm{Cc}_{b}({\mathcal{R}}(\sigma,V(\mathrm{R}))}b_{i}(C)\leq b_{i}(\overline{V}(\mathrm{R}))+b^{\dim\overline{V}-i-1}(\bigcup_{Q\in\mathcal{Q}}{\rm Z}(Q^{h},\overline{V}(\mathrm{R})).$
Proof.
Let $K=\bigcup_{Q\in\mathcal{Q}}{\rm Z}(Q^{h},\overline{V}(\mathrm{R}))$. Since $\overline{V}(\mathrm{R})$ is a compact (real) manifold (orientable over $\mathbb{Z}/2\mathbb{Z}$, and $K$ is a closed and bounded semi-algebraic subset, we obtain from Lefschetz duality theorem (see for example [43, page 297, Theorem 18]),  [^4] By a standard application of the Tarski-Seidenberg transfer principle we can assume that it holds over a general real closed field $\mathrm{R}$. that for each $i$, $\mbox{\rm H}_{i}(\overline{V}(\mathrm{R})-K)\cong\mbox{\rm H}^{\dim\overline{V}-i}(\overline{V}(\mathrm{R}),K)$. Moreover, using the cohomology long exact sequence one has
$\dim\mbox{\rm H}^{\dim\overline{V}-i}(\overline{V}(\mathrm{R}),K)\leq\dim\mbox{\rm H}^{\dim\overline{V}-i}(\overline{V}(\mathrm{R}))+\dim\mbox{\rm H}^{\dim\overline{V}-i-1}(K).$
Thus, we get
(4.3) $b_{i}(\overline{V}(\mathrm{R})-K)\leq b_{\dim\overline{V}-i}(\overline{V}(\mathrm{R}))+b_{\dim\overline{V}-i-1}(K).$
Now suppose that $C\in\mathrm{Cc}_{b}({\mathcal{R}}(\sigma,V(\mathrm{R}))$ for some $\sigma\in\{1,-1\}^{\mathcal{Q}}$. Then, $C\in\mathrm{Cc}(\overline{V}(\mathrm{R})-\bigcup_{Q\in\mathcal{Q}}{\rm Z}(Q^{h},\overline{V}(\mathrm{R})))$.
This implies that for each $i\geq 0$,
(4.4) $\sum_{\sigma\in\{1,-1\}^{\mathcal{Q}}}\sum_{C\in\mathrm{Cc}_{b}({\mathcal{R}}(\sigma,V(\mathrm{R}))}b_{i}(C)\leq b_{i}(\overline{V}(\mathrm{R})-\bigcup_{Q\in\mathcal{Q}}{\rm Z}(Q^{h},\overline{V}(\mathrm{R}))).$
The lemma now follows from inequalities (4.3) and (4.4). ∎
Lemma 4.17.
Let $\overline{V}\subset\mathbb{P}^{N}_{\mathrm{C}}$ be a real variety and $\mathcal{Q}=\{Q_{1},\ldots,Q_{s}\},Q_{i}\in\mathrm{R}[X_{0},\ldots,X_{N}]$ be a finite set of homogeneous polynomials. Then, for each $n\geq 0$,
$b_{n}(\overline{V}(\mathrm{R})\cap\bigcup_{i\in[1,s]}{\rm Z}(Q_{i},\mathbb{P}^{N}_{\mathrm{R}}))$$\leq$ $\sum_{i=0}^{n}\sum_{H\subset[1,s],\mathrm{card}(H)=i+1}b_{n-i}(\overline{V}(\mathrm{R})\cap\bigcap_{h\in H}{\rm Z}(Q_{h},\mathbb{P}^{N}_{\mathrm{R}}))$
$\leq$ $\sum_{i=0}^{n}\sum_{H\subset[1,s],\mathrm{card}(H)=i+1}b(\overline{V}\cap\bigcap_{h\in H}{\rm Z}(Q_{h},\mathbb{P}^{N}_{\mathrm{C}})).$
Proof.
The first inequality follows from the Mayer-Vietoris spectral sequence associated to the cover of the closed semi-algebraic set $S=\overline{V}(\mathrm{R})\cap\bigcup_{i\in I}{\rm Z}(Q_{i},\mathbb{P}^{N}_{\mathrm{R}})$ by the closed semi-algebraic subsets $S_{i}=\overline{V}(\mathrm{R})\cap{\rm Z}(Q_{i},\mathbb{P}^{N}_{\mathrm{R}})$. The $E_{1}$-page of this spectral sequence is given by
$E_{1}^{i,j}=\bigoplus_{H\subset[1,s],\mathrm{card}(H)=i+1}\mbox{\rm H}^{j}(\overline{V}(\mathrm{R})\cap\bigcap_{h\in H}{\rm Z}(Q_{h},\mathbb{P}^{N}_{\mathrm{R}})),$
with
$\mbox{\rm H}^{n}(S)=\bigoplus_{i+j=n}E_{\infty}^{i,j}.$
The first inequality in the lemma is an immediate consequence. ∎
Lemma 4.18 (Thom-Smith inequality).
Let $\overline{V}\subset\mathbb{P}^{N}_{\mathrm{C}}$ be a real variety and $Q\in\mathrm{R}[X_{0},\ldots,X_{N}]$ a homogeneous polynomial. Then,
$b(\overline{V}(\mathrm{R})\cap{\rm Z}(Q,\mathbb{P}^{N}_{\mathrm{R}}))\leq b(\overline{V}\cap{\rm Z}(Q,\mathbb{P}^{N}_{\mathrm{C}})).$
Proof.
If $X$ is a compact space (or a regular complex) equipped with an involution map $c:X\rightarrow S$, and $\mathrm{Fix}(c)\subset X$ denotes the subspace of fixed points of $X$. The Smith exact sequence (see for example [14, page 126]) then implies that
(4.5) $b(\mathrm{Fix}(c),\mathbb{Z}/2\mathbb{Z})$$\leq$ $b(X,\mathbb{Z}/2\mathbb{Z}).$
To prove the lemma, take the involution $c$ to be the complex conjugation. ∎
Lemma 4.19.
Let $\overline{V}\subset\mathbb{P}^{N}_{\mathrm{C}}$ be a complete intersection non-singular variety with $\dim\overline{V}=p,\deg(\overline{V})=D$. Let $P_{1},\ldots,P_{q}\in\mathrm{C}[X_{0},\ldots,X_{N}]$ be generic homogeneous polynomials with $\deg(P_{i})\leq d$, $D\leq d$, $q\leq p$. Then,
$b(\overline{V}\cap\bigcap_{i}{\rm Z}(P_{i},\mathbb{P}^{N}_{\mathrm{C}}))\leq 2^{N+1}\cdot D\cdot d^{p}.$
Proof.
If $\overline{V}\subset\mathbb{P}^{N}_{\mathrm{C}}$ is a non-singular complete intersection variety with $\dim\overline{V}=N-\ell$, and defined as the intersection of non-singular hypersurfaces of degrees $e_{1},\ldots,e_{\ell}$, then $b(\overline{V})$ is given by the following classical formula (see for example [16, §5.7.2-3])
$b(\overline{V})$ $=$ $(1+(-1)^{N-\ell+1})\cdot(N-\ell+1)+$
$e_{1}\cdot e_{2}\cdots e_{\ell}\cdot\left(\sum_{i=0}^{k-\ell}(-1)^{i}\cdot\binom{N+1}{i}\cdot h_{N-\ell-i}(e_{1},\ldots,e_{\ell})\right)$
$=$ $(1+(-1)^{N-\ell+1})\cdot(N-\ell+1)+$
$e_{1}\cdot e_{2}\cdots e_{\ell}\cdot\left(\sum_{j=0}^{N-\ell}(-1)^{N-\ell-j}\cdot\binom{N+1}{j+\ell+1}\cdot h_{j}(e_{1},\ldots,e_{\ell})\right),$
where $h_{j}(e_{1},\ldots,e_{\ell})$ is the complete homogeneous symmetric polynomial of degree $j$ in $(e_{1},\ldots,e_{\ell})$.
In our case, we can assume that the degree sequence of the polynomials defining $\overline{V}$ is of length $\ell=N-p+q$. Suppose that the degree sequence is $e_{1},\ldots,e_{N-p+q}$. We can assume without loss of generality that $e_{1}\cdot e_{N-p}=D$, and $e_{N-p+1}\leq d,\ldots,e_{N-p+q}\leq d$.
Thus,
(4.7) $e_{1}\cdots e_{\ell}\leq D\cdot d^{q},$
and
$\left(\sum_{j=0}^{N-\ell}(-1)^{N-\ell-j}\cdot\binom{N+1}{j+\ell+1}\cdot h_{j}(e_{1},\ldots,e_{\ell})\right)$$\leq$ $2^{N+1}h_{N-\ell}(e_{1},\ldots,e_{\ell})$
$\leq$ $2^{N+1}d^{N-\ell}$
$=$ $2^{N+1}d^{q}.$
The lemma now follows from the equality (4.3.2), and the inequalities (4.7) and (4.3.2). ∎
Proof of Theorem 5.
Let $\overline{V}\subset\mathbb{P}_{\mathrm{C}}^{N}$ be a non-singular complete intersection real variety with $\deg(V)=D$ and $\dim(V)=p$, and $V=\overline{V}\cap\mathrm{C}^{N}$.
Let $\mathcal{P}\subset\mathrm{R}[X_{1},\ldots,X_{N}]_{\leq d}$ be a finite set of polynomials with $d\geq D$ and $\mathrm{card}(\mathcal{P})=s$.
First observe that it follows from the theorem of conic structure at infinity [7, Proposition 5.49], that for all large enough $R>0$,
$\sum_{\sigma\in\{0,1,-1\}^{\mathcal{P}}}b_{0}({\mathcal{R}}(\sigma,V(\mathrm{R})))\leq\sum_{\sigma\in\{0,1,-1\}^{\mathcal{P}}}b_{0}({\mathcal{R}}(\sigma,V(\mathrm{R})\cap\overline{B_{N}(0,R)})).$
Now let $\mathcal{P}_{0}=\mathcal{P}\cup\{\sum_{i=1}^{N}X_{i}^{2}-R\}$. We have that
$\sum_{\sigma\in\{0,1,-1\}^{\mathcal{P}}}b_{i}({\mathcal{R}}(\sigma,V(\mathrm{R})))$$\leq$ $\sum_{\sigma\in\{0,1,-1\}^{\mathcal{P}}}b_{i}({\mathcal{R}}(\sigma,V(\mathrm{R})\cap\overline{B_{N}(0,R)}))$
$\leq$ $\sum_{\sigma^{\prime}\in\{0,1,-1\}^{\mathcal{P}_{0}},C\in\mathrm{Cc}_{b}({\mathcal{R}}(\sigma^{\prime},V(\mathrm{R})))}b_{i}(C)$
[…]
$\leq$ $b(\overline{V}(\mathrm{R}))+b_{p-i-1}(\bigcup_{\widetilde{P}\in\widetilde{\mathcal{P}}}{\rm Z}(\widetilde{P}^{h},\overline{V}(\mathrm{R}{\langle}{\varepsilon},\delta{\rangle})))$
[…]
The theorem now follows from Lemmas 4.17, 4.18 and 4.19. ∎
4.3.3. Outline of the proof of Theorem 6
In Theorem 6 (unlike in Theorem 5) we do not assume that the given complete intersection is non-singular. Our conclusion is also weaker, both from the point of view of the bound and the quantity being bounded.
The main idea is that after making an initial perturbation we can restrict our attention to strict conditions on the perturbed family of polynomials $\widetilde{P}$ (Lemma 4.20).
We now make an infinitesimal perturbation of the polynomials defining the complete intersection $V$ to obtain a non-singular complete intersection $\widetilde{V}$. (Lemma 4.21) We cannot ensure that if a semi-algebraically connected component $C$ of $\mathcal{P}$ meets $V(\mathrm{R})$, then the corresponding semi-algebraically connected component $\widetilde{C}$ of the realization of a strict sign condition on $\widetilde{P}$ will meet $\widetilde{V}(\mathrm{R})$. Nevertheless, identifying $\mathrm{C}^{N}$ with $\mathrm{R}^{2N}$, and the interpreting the polynomials in $\mathcal{P}$ and $\widetilde{\mathcal{P}}$ as polynomials in $2N$ variables, it is true that the variety $\widetilde{V}\subset\mathrm{C}^{N}=\mathrm{R}^{2N}$ will meet $\widetilde{C}\times\widetilde{\mathrm{R}}^{N}$ (Lemma 4.23).
Moreover, separating the real and complex parts of the polynomials defining the variety $\widetilde{V}$ one obtains a complex non-singular complete intersection in $\mathbb{P}^{2N}_{\mathrm{C}}$, of degree $D^{2}$ and dimension $2p$. We can now apply Theorem 5.
4.3.4. Proof of Theorem 6
We need the following lemmas.
Lemma 4.20.
Let $\mathcal{P}=\{P_{1},\ldots,P_{s}\}\subset\mathrm{R}[X_{1},Y_{1},\ldots,X_{N},Y_{N}]$ be a finite set, and $P_{0}=\sum_{i=1}^{N}(X_{i}^{2}+Y_{i}^{2})-R$, where $R>0,R\in\mathrm{R}$. For each $i,0\leq i\leq s$, let $H_{i}\in\mathrm{R}[X_{1},Y_{1},\ldots,X_{N},Y_{N}]$ be a generic polynomial with, $H_{i}(x)>0$ for all $x\in\overline{B_{2N}(0,R)}$, and $\deg(H_{i})=\deg(P_{i})$.
For each $\sigma\in\{0,1,-1\}^{\mathcal{P}}$, and $C\in\mathrm{Cc}({\mathcal{R}}(\sigma,\overline{B_{2N}(0,R)})$, there exists a unique $\widetilde{\sigma}\in\{-1,1\}^{\widetilde{\mathcal{P}}}$ and a unique semi-algebraically connected component of $\widetilde{C}$ of ${\mathcal{R}}(\widetilde{\sigma},\overline{B_{2N}(0,R)})$, such that $\widetilde{C}\cap\mathrm{R}^{2N}=C$, where
$\widetilde{P}=\bigcup_{0\leq i\leq s}\{P_{i}\pm\delta_{i}H_{i},P_{i}\pm\eta_{i}H_{i}\}.$
Proof.
Similar to the proof of Lemma 4.14 and omitted. ∎
Lemma 4.21.
Let $V\subset\mathrm{C}^{N}$ be a complete intersection variety of dimension $p$ defined by polynomials $\mathcal{Q}=\{Q_{1},\ldots,Q_{N-p}\}\subset\mathrm{C}[Z_{1},\ldots,Z_{N}]$.
Then there exists polynomials $\widetilde{\mathcal{Q}}=\widetilde{Q}_{1},\ldots,\widetilde{Q}_{N-p}\subset\mathrm{C}{\langle}{\varepsilon}{\rangle}[Z_{1},\ldots,Z_{N}]$ such that:
  1. (A)
    for each $i,1\leq i\leq N-p$, $\deg(\widetilde{Q}_{i})=\deg(Q_{i})$;
  2. (B)
    the homogenizations $\widetilde{Q}_{1}^{h},\ldots,\widetilde{Q}_{N-p}^{h}$ define a non-singular complete intersection in $\mathbb{P}_{\mathrm{C}{\langle}{\varepsilon}{\rangle}}^{N}$;
  3. (C)
    for all $R\in\mathrm{R},R>0$, $\lim_{\varepsilon}{\rm Z}(\widetilde{\mathcal{Q}},B_{N}(0,R+1))\cap\overline{B_{N}(0,R)}={\rm Z}(\mathcal{Q},\overline{B(0,R)})$, where $B(0,R)\subset\mathrm{C}^{N}$ is the open ball of radius $R$ in $\mathrm{C}^{N}=\mathrm{R}^{2N}$.
Proof.
For $1\leq i\leq N-p$, let $H_{1},\ldots,H_{N-p}$ be generically chosen polynomials with $\deg(H_{i})=\deg(Q_{i})$. Now let $\widetilde{Q}_{i}=(1-{\varepsilon})Q_{i}+{\varepsilon}H_{i}$. Then properties (A) and (B) follows from the fact that same holds for the polynomials $H_{1},\ldots,H_{N-p}$ and a standard transfer argument.
For property (C) first observe that clearly $\lim_{\varepsilon}{\rm Z}(\widetilde{\mathcal{Q}},B(0,R+1))\subset{\rm Z}(\mathcal{Q},\overline{B_{N}(0,R+1)})$ from which it follows that
$\lim_{\varepsilon}{\rm Z}(\widetilde{\mathcal{Q}},B_{N}(0,R+1))\cap\overline{B_{N}(0,R)}={\rm Z}(\mathcal{Q},\overline{B_{N}(0,R)}).$
To prove the reverse inclusion let $x\in(\mathcal{Q},\overline{B(0,R)})$. We prove that there exists $\widetilde{x}\in{\rm Z}(\widetilde{\mathcal{Q}},B_{N}(0,R+1))$ such that $\lim_{\varepsilon}\widetilde{x}=x$.
Let $\mathbb{P}^{N-p}_{\mathrm{C}}\subset\mathbb{P}^{N}_{\mathrm{C}}$ be a generic $(N-p)$-dimensional plane through $x$. Then, using Bezout’s theorem ${\rm Z}(\mathcal{Q}^{h},\mathbb{P}^{N-p}_{\mathrm{C}})$ contains $x$ and has $d_{1}\cdots d_{N-p}$ points counted with multiplicities. But ${\rm Z}(\widetilde{\mathcal{Q}}^{h},\mathbb{P}^{N-p}_{\mathrm{C}})$ also contains $d_{1}\cdots d_{N-p}$ points each of multiplicity one, and moreover $\lim_{\varepsilon}{\rm Z}(\widetilde{\mathcal{Q}}^{h},\mathbb{P}^{N-p}_{\mathrm{C}})={\rm Z}(\widetilde{\mathcal{Q}}^{h},\mathbb{P}^{N-p}_{\mathrm{C}})$. From this it follows that there must exists $\widetilde{x}\in{\rm Z}(\widetilde{\mathcal{Q}}^{h},\mathbb{P}^{N-p}_{\mathrm{C}})$ such that $\lim_{\varepsilon}\widetilde{x}=x$, and clearly $\widetilde{x}\in{\rm Z}(\widetilde{\mathcal{Q}},B_{N}(0,R+1))$. ∎
Now let $V\subset\mathrm{C}^{N}$ be a complete intersection variety of dimension $p$ defined by polynomials $\mathcal{Q}=\{Q_{1},\ldots,Q_{N-p}\}\subset\mathrm{C}[Z_{1},\ldots,Z_{N}]$, and let $\widetilde{\mathcal{Q}}$ be the set of corresponding polynomials in Lemma 4.20.
Lemma 4.22.
Let $\widetilde{P}\subset\mathrm{R}[X_{1},Y_{1},\ldots,X_{N},Y_{N}]$ be a finite set, $\sigma\in\{-1,1\}^{\widetilde{\mathcal{P}}}$, $C\in\mathrm{Cc}({\mathcal{R}}(\sigma,\mathrm{C}^{N}))$ such that $C\cap V\neq\emptyset$. Then, ${\rm Ext}(C,\mathrm{C}{\langle}{\varepsilon}{\rangle})\cap\widetilde{V}\neq\emptyset$, where $\widetilde{V}={\rm Z}(\widetilde{\mathcal{Q}},\mathrm{C}{\langle}{\varepsilon}{\rangle}^{N})$.
Proof.
This follows immediately from Lemma 4.20 noting that ${\mathcal{R}}(\sigma,\mathrm{C}^{N})$ is an open semi-algebraic subset. ∎
Lemma 4.23.
Let $V,\widetilde{V}$ be as in Lemma 4.21 and $\mathcal{P},\widetilde{\mathcal{P}}$ be as in Lemma 4.20. Then,
$\sum_{\sigma\in\{0,1,-1\}^{\mathcal{P}}}\mathrm{card}(\mathrm{Cc}({\mathcal{R}}(\sigma),V(\mathrm{R})))\leq\sum_{\widetilde{\sigma}\in\{1,-1\}^{\widetilde{\mathcal{P}}}}\mathrm{card}(\mathrm{Cc}({\mathcal{R}}(\widetilde{\sigma}),\widetilde{V}))\leq\sum_{\widetilde{\sigma}\in\{1,-1\}^{\widetilde{\mathcal{P}}}}b_{0}({\mathcal{R}}(\widetilde{\sigma},\widetilde{V})).$
Proof.
Follows immediately from (Lemma 4.22). ∎
Proof of Theorem 6.
Using Lemma 4.23, it suffices to bound
$\sum_{\widetilde{\sigma}\in\{1,-1\}^{\widetilde{\mathcal{P}}}}b_{0}({\mathcal{R}}(\widetilde{\sigma},\widetilde{V})).$
Now let for $j=1,\ldots,N-p$, let
$\widetilde{Q}_{j}=\widetilde{Q}_{j,0}+\sqrt{-1}\cdot\widetilde{Q}_{j,1},$
where $\widetilde{Q}_{j,0},\widetilde{Q}_{j,1}\in\mathrm{R}[X_{1},Y_{1},\ldots,X_{n},Y_{n}]$ have generic coefficients.
Thus,
$\widetilde{Q}_{1,0},\widetilde{Q}_{1,1},\ldots,\widetilde{Q}_{N-p,0},\widetilde{Q}_{N-p,1}$
define a non-singular complete intersection in $\widetilde{W}\subset\mathrm{C}^{2N}$, with $\deg(\widetilde{W})=D^{2}$, and $\dim\widetilde{W}=2p$ and with $\widetilde{W}(\mathrm{R})=\widetilde{V}$.
Now apply Theorem 5 with non-singular complete intersection $\widetilde{W}$ and the family of polynomials $\widetilde{\mathcal{P}}$ to obtain
$\sum_{\widetilde{\sigma}\in\{1,-1\}^{\widetilde{\mathcal{P}}}}b_{0}({\mathcal{R}}(\widetilde{\sigma},\widetilde{V}))$$\leq$ $\sum_{\sigma\in\{0,1,-1\}^{\widetilde{\mathcal{P}}}}b_{0}({\mathcal{R}}(\widetilde{\sigma},\widetilde{W}(\mathrm{R})))$
$\leq$ $(D^{2})^{2}\cdot(O(sd))^{2p}$
$=$ $D^{4}\cdot(O(sd))^{2p}.$
4.4. Proof of Theorem 7
4.4.1. Outline of the proof of Theorem 7
The proof borrows from the proof of Theorem 5 the main technique of making perturbations of the given family $\mathcal{P}$, and using the inequalities arising from the Mayer-Vietoris spectral sequence to reduce to the case of bounding the Betti numbers of certain non-singular projective varieties. Instead of using the Thom-Smith inqualities and the Chern class formula as we do in the proof of Theorem 5, we make a further reduction using the same technique as in the proof of Theorem 2 to reduce to bounding the Betti numbers of non-singular real projective varieties of dimensions $p-1$ (and also bounding that of $V(\mathrm{R})$ which is of dimension $p$). We use in this case the bound provided by Theorem F and the adjoining Remark 4 to bound these.
4.4.2. Proof of Theorem 7
We already have the necessary ingredients to start the proof without having to proof additional lemmas.
Proof of Theorem 7.
Using the conic structure at infinity for semi-algebraic sets (see for example [7, Proposition 5.49]) choose $R>0,R\in\mathrm{R}$, such that for each $\sigma\in\{0,1,-1\}^{\mathcal{P}}$, ${\mathcal{R}}(\sigma,V(\mathrm{R}))$ is semi-algebraically homeomorphic to ${\mathcal{R}}(\sigma,V(\mathrm{R})\cap B_{N}(0,R))$. We denote $\mathcal{P}_{0}=\mathcal{P}\cup P_{0}$, where $P_{0}=\sum_{i=1}^{N}X_{i}^{2}-R$.
Applying Lemma 4.15 (using the same notation as in Lemma 4.15) we have
[…]
and finally using Lemma 4.16 that,
[…]
We now use the Mayer-Vietoris spectral sequence to observe that the right hand side of the previous equation is bounded by the sum of three terms (we use the convenient notation $A\subset_{n}B$ to denote $A\subset B$ and $\mathrm{card}(A)=n$):
  1. (a)
    $\sum_{0\leq j\leq p-i-1}\sum_{\widetilde{\mathcal{P}}_{0}^{\prime}\subset_{j+1}\widetilde{\mathcal{P}}_{0}}b_{p-i-1-j}({\rm Z}((\widetilde{\mathcal{P}}_{0}^{\prime})^{h},{\rm Ext}(\overline{V}(\mathrm{R}),\widetilde{\mathrm{R}}))),$
  2. (b)
    $\sum_{0\leq j\leq p-i-2}\sum_{\widetilde{\mathcal{P}}_{0}^{\prime}\subset_{j+1}\widetilde{\mathcal{P}}_{0}}b_{p-i-1-j}({\rm Z}(\{X_{0}\}\cup(\widetilde{\mathcal{P}}_{0}^{\prime})^{h},{\rm Ext}(\overline{V}(\mathrm{R}),\widetilde{\mathrm{R}}))),$
  3. (c)
    $b({\rm Ext}(\overline{V}(\mathrm{R}),\widetilde{\mathrm{R}})).$
We are now going to bound from above each one of the three items. Firstly, observe that using Remark 4 and the fact that $V$ is non-singular we can bound the quantity in Part (c) by $D^{p+1}$.
We now bound for each $\widetilde{\mathcal{P}}_{0}^{\prime}\subset_{j+1}\widetilde{\mathcal{P}}_{0}$, the Betti number $b_{p-i-1-j}({\rm Z}((\widetilde{\mathcal{P}}_{0}^{\prime})^{h},{\rm Ext}(\overline{V}(R),\widetilde{\mathrm{R}})))$.
Note that ${\rm Z}((\widetilde{\mathcal{P}}_{0}^{\prime})^{h},{\rm Ext}(\overline{V}(\mathrm{R}),\widetilde{\mathrm{R}}))$ is the real part of a non-singular projective hypersurface with degree bounded by $Dd^{j+1}$ and dimension $p-j-1$. Using Remark 4 and the Thom-Smith inequality will give us
$b_{p-i-1-j}({\rm Z}((\widetilde{\mathcal{P}}_{0}^{\prime})^{h},{\rm Ext}(\overline{V}(\mathrm{R}),\widetilde{\mathrm{R}})))\leq(Dd^{j+1})^{p-j}=D^{p-j}d^{(j+1)(p-j)}.$
Using perturbation techniques we prove a slightly better bound as follows.
Let $S={\rm Z}((\widetilde{\mathcal{P}}_{0}^{\prime})^{h},{\rm Ext}(\overline{V}(\mathrm{R}),\widetilde{\mathrm{R}}))$.
First let
$\widetilde{Q}=\sum_{\widetilde{P}\in\widetilde{\mathcal{P}}_{0}^{\prime}}\widetilde{P}^{2}.$
Let $H\in\widetilde{\mathrm{R}}[X_{0},\ldots,X_{N}]$ be a generic homogeneous polynomial of the same degree as $\widetilde{Q}$ and which in non-negative in $\mathrm{R}^{N+1}$. Note that in $\mathbb{P}_{\mathrm{R}}^{N}$ the sign of a homogeneous polynomial of even degree is well defined. It follows from [7, Lemma 16.17] that the semi-algebraic set $\widetilde{S}\subset{\rm Ext}(\overline{V}(R),\widetilde{R}{\langle}\zeta{\rangle}))$ defined as the intersection of ${\rm Ext}(\overline{V}(R),\widetilde{R}{\langle}\zeta{\rangle}))$ with the semi-algebraic set defined by $(1-\zeta)\widetilde{Q}-\zeta H\leq 0$ is semi-algebraically homotopy equivalent to ${\rm Ext}(S,\widetilde{\mathrm{R}}{\langle}\zeta{\rangle})$.
Moreover, it follows from the Mayer-Vietoris exact sequence that
$b(\widetilde{S})\leq b({\rm Z}((1-\zeta)\widetilde{Q}-\zeta H,{\rm Ext}(\overline{V}(\mathrm{R}),\widetilde{\mathrm{R}}{\langle}\zeta{\rangle})))+b(V(\mathrm{R})).$
Now observe that ${\rm Z}((1-\zeta)\widetilde{Q}-\zeta H,{\rm Ext}(\overline{V}(\mathrm{R}),\widetilde{\mathrm{R}}{\langle}\zeta{\rangle}))$ is the real part of a non-singular projective variety of degree bounded by $2Dd$ and dimension $p-1$. It follows (using Remark 4 and the Thom-Smith inequality again) that
$b({\rm Z}((1-\zeta)\widetilde{Q}-\zeta H,{\rm Ext}(\overline{V}(\mathrm{R}),\widetilde{\mathrm{R}}{\langle}\zeta{\rangle})))\leq(2Dd)^{p}.$
Thus, the sum in Part (a) is bounded by
$\left(\sum_{j=0}^{p-i-1}\binom{s}{j+1}\right)(D^{p}(2d)^{p}+D).$
We use the same method to bound the sum in Part (b), noting that all subvarieties are contained in the $\mathbb{P}_{\mathrm{R}}^{N-1}$ defined by $X_{0}=0$. This yields that the sum in Part (b) is bounded by
$\left(\sum_{j=0}^{p-i-2}\binom{s}{j+1}\right)(D^{p-1}(2d)^{p-1}+D).$
Finally, summing the three parts together we obtain that:
$\sum_{\sigma\in\{0,1-1\}^{\mathcal{P}}}b_{i}({\mathcal{R}}(\sigma,V(\mathrm{R})))$$\leq$ $\left(\sum_{j=0}^{p-i-1}\binom{s}{j+1}\right)(D^{p}(2d)^{p}+D)+$
$\left(\sum_{j=0}^{p-i-2}\binom{s}{j+1}\right)(D^{p-1}(2d)^{p-1}+D)+$
$D^{p+1}$
$\leq$ $O(1)^{p}s^{p-i}D^{p}(d^{p}+D).$
4.5. Proofs of Theorems 8 and 9
Proof of Theorem 8.
We only consider the affine case. The projective case can be reduced to the affine case using by taking affine covers. For each $n>0$, let
$\mathcal{Q}_{n}\subset k[\mathbf{X}^{(1)},\ldots,\mathbf{X}^{(n)}],$
be defined by
$\mathcal{Q}_{n}=\bigcup_{1\leq i,j\leq n}\bigcup_{P\in\mathcal{P}_{n}}P(\mathbf{X}^{(i)},\mathbf{X}^{(j)}),$
where each $\mathbf{X}^{(i)}$ is a block of $N_{n}$-variables.
Then, then it is easy to verify that the of number distinct $(Z_{n},F_{n})$-graphs is bounded by $\mathrm{card}(\Pi(\mathcal{Q}_{n},Z_{n}^{n}))$.
Now, $\deg(Z_{n}^{n})=\deg(Z_{n})^{n}\leq c_{3}^{n}n^{n}$, $\dim Z_{n}^{n}\leq np$, $\mathrm{card}(\mathcal{Q}_{n})=c_{1}n^{2}$, $\deg(Q)\leq c_{2}n$ for $Q\in\mathcal{Q}_{n}$.
Applying Theorem 1 we obtain
$\mathrm{card}(\Pi(\mathcal{Q}_{n},Z_{n}^{n}))$$\leq$ $c_{3}^{n}n^{n}\sum_{i=0}^{np}\binom{c_{1}n^{2}}{i}(c_{2}n)^{pn-i}$
$\leq$ $c_{3}^{n}n^{n}\sum_{i=0}^{np}\left(\frac{ec_{1}n^{2}}{i}\right)^{i}(c_{2}n)^{pn-i}$
$=$ $c_{3}^{n}\cdot n^{n}\cdot\sum_{i=0}^{np}(ec_{1})^{i}\cdot(c_{2})^{pn-i}\cdot n^{pn}\left(\frac{n}{i}\right)^{i}$
$\leq$ $n^{(1+o(1))(p+1)n}.$
Proof of Theorem 9.
The proof is very similar to the proof of Theorem 8 and we omit the details. In the last step, we use Theorem 6 instead of Theorem 1 to bound the number of realizable sign conditions. The bound that we obtain is slightly worse (in the exponent) than the bound in Theorem 8 reflecting the fact that the bound on the realizable sign conditions in Theorem 6 is a bit worse than the one in Theorem 1. ∎
4.6. Proof of Theorem 10
We first recall a notion from [51] that plays a key role.
Definition 4.1.
A set $V\subset\mathrm{R}^{N}$ is called a $(K,n)$-regular set if
  1. (a)
    For almost all affine planes $L$ of codimension $n^{\prime}\leq n$ in $\mathrm{R}^{N}$, $V\cap L$ has at most $K$ path connected components.
  2. (b)
    For almost all affine planes $L$ of codimension $n^{\prime}\geq n+1$ in $\mathrm{R}^{N}$, $V\cap L$ is empty.
The main theorem in [51] is the following theorem.
Theorem J.
Let $V$ be $(K,n)$ regular set in the cube $[-1,1]^{N}$, the covering number is bounded by
(4.9) $\log\mathcal{N}(V,\varepsilon)\leq n\log(1/\varepsilon)+\log K+Cn\log N,$
for some absolute constant $C$.
Proof of Theorem 10.
The proof is now an immediate corollary of inequality (4.9) and inequality (2.1) in Theorem 3. ∎
4.7. Proof of Theorem 11
Proof of Theorem 11.
First suppose that $S\subset V(\mathrm{R})$, and $V$ is a union of non-singular real varieties, with $\dim V=p$ and $\deg V=D$. Suppose there exists an algebraic computation tree of height $t$ testing membership in $S$. Then, the number of leaves of the tree (in particular the number of leaves labelled YES) is bounded by $2^{t}$. Let $w$ denote a leaf of the tree labelled YES. Then the path from the root to $w$ defines a basic semi-algebraic set $S_{w}\subset V(\mathrm{R})\times\mathrm{R}^{t_{w}}$, where $t_{w}\leq t$ is the number of variables introduced in the path to $w$ (i.e. number of computation vertices on this path, see Definition 3.3), and $S_{w}$ is defined by a set $\mathcal{P}_{w}$ of at most $t$ polynomials of degree at most $2$.
Then, $b_{0}(S_{w})$ is bounded by $2^{t}O(D)^{p+t_{w}}=O(D)^{p+t}$ using Theorem 2.
This implies that
$b_{0}(S)\leq O(D)^{p+t}.$
Taking logarithm of both sides we obtain the desired lower bound on $t$. ∎
4.8. Proofs of Theorems 12 and 13
Proof of Theorem 12.
If $\mathrm{rank}_{A,\Delta}(W)=\infty$, then there is nothing to prove. Otherwise, let $\mathrm{rank}_{A,\Delta}(W)=t_{0}$.
For $1\leq t\leq t_{0}$, let $V_{t}=A^{t}$, and let $V=\oplus_{1\leq t\leq t_{0}}V_{t}$ denote the direct sum and $i_{t}:V_{t}\rightarrow V$ the canonical injection. Let $\Delta^{\prime}=\cup_{1\leq t\leq t_{0}}i_{t}(\Sigma^{t})$, noting that $\Delta^{t}\subset V_{t}$ is an algebraic subset of $V_{t}$.
Notice that $\Delta^{\prime}$ is an algebraic subset of dimension bounded by $pt_{0}$ and degree bounded by $D+D^{2}+\cdots+D^{t_{0}}\leq D^{t_{0}+1}$. Finally, let $\phi:V\rightarrow A$ denote the map induced (by extending linearly) the maps $\phi_{t}:V_{t}\rightarrow A$, sending $(v_{1},\ldots,v_{t})$ to $v_{1}+\cdots+v_{t}$ (i.e. $\phi=\sum_{1\leq t\leq t_{0}}\phi_{t}\circ\pi_{t}$, where $\pi_{t}:V\rightarrow V_{t}$ is the canonically defined projection).
Then, $\phi_{t}^{-1}(W)\cap\Delta^{t}$ is an algebraic subset of $\Delta^{t}$ cut out by at most $s$ polynomials of degrees bounded by $\delta$, and so $\phi^{-1}(W)\cap\Delta^{\prime}$ is an algebraic subset of $\Delta^{\prime}$ cut out by at most $st_{0}$ polynomials of degrees bounded by $\delta$.
Since for $u,v\in V,u\neq v\neq 0$, $\phi^{-1}(u)\cap\Delta^{\prime}$ and $\phi^{-1}(v)\cap\Delta^{\prime}$ are disjoint,
$b_{0}((\phi^{-1}(u)\cup\phi^{-1}(v))\cap\Delta^{\prime})=b_{0}(\phi^{-1}(u)\cap\Delta^{\prime})+b_{0}(\phi^{-1}(v)\cap\Delta^{\prime})$.
More generally,
$b_{0}(\phi^{-1}(W)\cap\Delta^{\prime})=\sum_{v\in W}b_{0}(\phi^{-1}(v)\cap\Delta^{\prime}).$
Moreover, since
$\{v\in W\;|\;\phi^{-1}(W)\cap\Delta^{\prime}\neq\emptyset\}=W$
we have
$b_{0}(\phi^{-1}(W)\cap\Delta^{\prime})$ $=$ $\sum_{v\in W}b_{0}(\phi^{-1}(v)\cap\Delta^{\prime})$
$\geq$ $\mathrm{card}(\{v\in W\;|\;\phi^{-1}(W)\cap\Delta^{\prime}\neq\emptyset\})$
$=$ $\mathrm{card}(W).$
It follows from Theorem 1 that
$\deg\overline{\phi^{-1}(W)\cap\Delta^{\prime})}$$\leq$ $b_{0}(\phi^{-1}(W)\cap\Delta^{\prime})$
$\leq$ $(pt_{0}+1)\cdot(t_{0}s)^{pt_{0}}D^{t_{0}+1}\cdot(\delta)^{pt_{0}}.$
Hence,
$\mathrm{card}(W)\leq(pt_{0}+1)\cdot(t_{0}s)^{pt_{0}}\cdot D^{t_{0}+1}\cdot(t_{0})^{pt_{0}}.$
Taking logarithm of both sides this implies,
$t_{0}\geq\frac{\log\mathrm{card}(W)}{O(p(\log s+\log\delta+\log\log\mathrm{card}(W))+\log D)},$
which finishes the proof. ∎
Proof of Theorem 13.
We follow the proof of Theorem 12. If $\mathrm{mmrank}_{V,\Delta}(W)=\infty$, then there is nothing to prove. Otherwise, let $\mathrm{mmrank}_{V,\Delta}(W)=t_{0}$.
For $1\leq t\leq t_{0}$, let $V^{*}=\oplus_{1\leq t\leq t_{0}}V^{t}$ denote the direct sum and $i_{t}:V^{t}\rightarrow V$ the canonical injection. Let $\Delta^{\prime}=\cup_{1\leq t\leq t_{0}}i_{t}(\Delta^{t})$, noting that $\Delta^{t}\subset UV^{t}L$ is a real algebraic subset of $V^{t}$.
Notice that $\Delta^{\prime}$ is a union of non-singular complete intersection subvarieties of dimension bounded by $pt_{0}$ and degree bounded by $D+D^{2}+\cdots+D^{t_{0}}\leq D^{t_{0}+1}$. Finally, as before let $\phi:V^{*}\rightarrow V$ denote the map induced (by extending linearly) the maps $\phi_{t}:V^{t}\rightarrow V$, sending $(v_{1},\ldots,v_{t})$ to $v_{1}+\cdots+v_{t}$ (i.e. $\phi=\sum_{1\leq t\leq t_{0}}\phi_{t}\circ\pi_{t}$, where $\pi_{t}:V^{*}\rightarrow V^{t}$ is the canonically defined projection).
Since for $\sigma,\sigma^{\prime}\in\{0,1,-1\}^{\mathcal{P}}$, $\phi^{-1}({\mathcal{R}}(\sigma))\cap\Delta^{\prime}$ and $\phi^{-1}({\mathcal{R}}(\sigma^{\prime}))\cap\Delta^{\prime}$ are disjoint,
$b_{0}((\phi^{-1}({\mathcal{R}}(\sigma))\cup\phi^{-1}({\mathcal{R}}(\sigma^{\prime}))\cap\Delta^{\prime})\leq b_{0}(\phi^{-1}({\mathcal{R}}(\sigma))\cap\Delta^{\prime})+b_{0}(\phi^{-1}({\mathcal{R}}(\sigma^{\prime}))\cap\Delta^{\prime}).$
We have
$\sum_{\sigma\in\{0,1,-1\}^{\mathcal{P}}}b_{0}((\phi^{-1}({\mathcal{R}}(\sigma))\cap\Delta^{\prime})$$\geq$ $b_{0}(W).$
It follows from Theorem 5 that
$\sum_{\sigma\in\{0,1,-1\}^{\mathcal{P}}}b_{0}(\phi^{-1}({\mathcal{R}}(\sigma))\cap\Delta^{\prime})\leq D^{2}(t_{0}s\delta)^{O(pt_{0})}.$
Hence,
$b_{0}(W)\leq D^{2}(t_{0}s\delta)^{O(pt_{0})}.$
Taking logarithm of both sides this implies,
$t_{0}\geq\frac{\log b_{0}(W)}{O(p(\log s+p\log\delta+\log\log b_{0}(W))+\log D)},$
which finishes the proof. ∎
4.9. Proofs of Theorems 14 and 15
Proof of Theorem 14.
If $\mathrm{rank}_{A,\Delta}(W)=\infty$, then there is nothing to prove. Otherwise, let $\mathrm{rank}_{A,\Delta}(W)=t_{0}$.
For $1\leq t\leq t_{0}$, let $V_{t}=A^{t}$, and let $V=\oplus_{1\leq t\leq t_{0}}V_{t}$ denote the direct sum and $i_{t}:V_{t}\rightarrow V$ the canonical injection. Let $\Delta^{\prime}=\cup_{1\leq t\leq t_{0}}i_{t}(\Sigma^{t})$, noting that $\Delta^{t}\subset V_{t}$ is an algebraic subset of $V_{t}$.
Notice that $\Delta^{\prime}$ is an algebraic subset of dimension bounded by $pt_{0}$ and degree bounded by $D+D^{2}+\cdots+D^{t_{0}}\leq D^{t_{0}+1}$. Finally, let $\phi:V\rightarrow A$ denote the map induced (by extending linearly) the maps $\phi_{t}:V_{t}\rightarrow A$, sending $(f_{1},\ldots,f_{t})$ to $f_{1}\cdots f_{t}$ (i.e. $\phi=\sum_{1\leq t\leq t_{0}}\phi_{t}\circ\pi_{t}$, where $\pi_{t}:V\rightarrow V_{t}$ is the canonically defined projection).
Then, $\phi_{t}^{-1}(W)\cap\Delta^{t}$ is an algebraic subset of $\Delta^{t}$ cut out by at most $s$ polynomials of degrees bounded by $t\delta$, and so $\phi^{-1}(W)\cap\Delta^{\prime}$ is an algebraic subset of $\Delta^{\prime}$ cut out by at most $st_{0}$ polynomials of degrees bounded by $t_{0}\delta$.
Since for $f,g\in A,f\neq g\neq 0$, $\phi^{-1}(f)\cap\Delta^{\prime}$ and $\phi^{-1}(g)\cap\Delta^{\prime}$ are disjoint,
$b_{0}((\phi^{-1}(f)\cup\phi^{-1}(g))\cap\Delta^{\prime})=b_{0}(\phi^{-1}(f)\cap\Delta^{\prime})+b_{0}(\phi^{-1}(g)\cap\Delta^{\prime})$.
More generally,
$b_{0}(\phi^{-1}(W)\cap\Delta^{\prime})=\sum_{f\in W}b_{0}(\phi^{-1}(f)\cap\Delta^{\prime}).$
Moreover, since
$\{f\in W\;|\;\phi^{-1}(W)\cap\Delta^{\prime}\neq\emptyset\}=W$
we have
$b_{0}(\phi^{-1}(W)\cap\Delta^{\prime})$ $=$ $\sum_{f\in W}b_{0}(\phi^{-1}(f)\cap\Delta^{\prime})$
$\geq$ $\mathrm{card}(\{f\in W\;|\;\phi^{-1}(W)\cap\Delta^{\prime}\neq\emptyset\})$
$=$ $\mathrm{card}(W).$
It follows from Theorem 1 that
$b_{0}(\phi^{-1}(W)\cap\Delta^{\prime})\leq(pt_{0}+1)\cdot(t_{0}s)^{pt_{0}}D^{t_{0}+1}\cdot(t_{0}\delta)^{pt_{0}}.$
Hence,
$\mathrm{card}(W)\leq(pt_{0}+1)\cdot(t_{0}s)^{pt_{0}}\cdot D^{t_{0}+1}\cdot(t_{0}\delta)^{pt_{0}}.$
Taking logarithm of both sides this implies,
$t_{0}\geq\frac{\log\mathrm{card}(W)}{O(p(\log s+\log\delta+\log\log\mathrm{card}(W))+\log D)},$
which finishes the proof. ∎
Proof of Theorem 15.
We follow the proof of Theorem 14. If $\mathrm{mmrank}_{A,\Delta}(W)=\infty$, then there is nothing to prove. Otherwise, let $\mathrm{mmrank}_{A,\Delta}(W)=t_{0}$.
For $1\leq t\leq t_{0}$, let $V_{t}=A^{t}$, and let $V=\oplus_{1\leq t\leq t_{0}}V_{t}$ denote the direct sum and $i_{t}:V_{t}\rightarrow V$ the canonical injection. Let $\Delta^{\prime}=\cup_{1\leq t\leq t_{0}}i_{t}(\Delta^{t})$, noting that $\Delta^{t}\subset V_{t}$ is a real algebraic subset of $V_{t}$.
Notice that $\Delta^{\prime}$ is a union of non-singular complete intersection subvarieties of dimension bounded by $pt_{0}$ and degree bounded by $D+D^{2}+\cdots+D^{t_{0}}\leq D^{t_{0}+1}$. Finally, as before let $\phi:V\rightarrow A$ denote the map induced (by extending linearly) the maps $\phi_{t}:V_{t}\rightarrow A$, sending $(f_{1},\ldots,f_{t})$ to $f_{1}\cdots f_{t}$ (i.e. $\phi=\sum_{1\leq t\leq t_{0}}\phi_{t}\circ\pi_{t}$, where $\pi_{t}:V\rightarrow V_{t}$ is the canonically defined projection).
Since for $\sigma,\sigma^{\prime}\in\{0,1,-1\}^{\mathcal{P}}$, $\phi^{-1}({\mathcal{R}}(\sigma))\cap\Delta^{\prime}$ and $\phi^{-1}({\mathcal{R}}(\sigma^{\prime}))\cap\Delta^{\prime}$ are disjoint,
$b_{0}((\phi^{-1}({\mathcal{R}}(\sigma))\cup\phi^{-1}({\mathcal{R}}(\sigma^{\prime}))\cap\Delta^{\prime})\leq b_{0}(\phi^{-1}({\mathcal{R}}(\sigma))\cap\Delta^{\prime})+b_{0}(\phi^{-1}({\mathcal{R}}(\sigma^{\prime}))\cap\Delta^{\prime}).$
We have
$\sum_{\sigma\in\{0,1,-1\}^{\mathcal{P}}}b_{0}((\phi^{-1}({\mathcal{R}}(\sigma))\cap\Delta^{\prime})$$\geq$ $b_{0}(W).$
It follows from Theorem 5 that
$\sum_{\sigma\in\{0,1,-1\}^{\mathcal{P}}}b_{0}(\phi^{-1}({\mathcal{R}}(\sigma))\cap\Delta^{\prime})\leq D^{2}(t_{0}s\delta)^{O(pt_{0})}.$
Hence,
$b_{0}(W)\leq D^{2}(t_{0}s\delta)^{O(pt_{0})}.$
Taking logarithm of both sides this implies,
$t_{0}\geq\frac{\log b_{0}(W)}{O(p(\log s+p\log\delta+\log\log b_{0}(W)+\log D)},$
which finishes the proof. ∎
4.10. Proofs of Theorems 16 and 17
Proof of Theorem 16.
Recall that for any Boolean function $f\in\{0,1\}^{n}\rightarrow\{0,1\}$, the stringent quantum circuit complexity of $f$ relative to $\Delta$ is by definition $\mathrm{rank}_{\mathrm{End}(V^{\otimes(n+1)},\Delta}(U_{f})$, where $V$ is a two-dimensional complex inner product space. Let $W=\{U_{f}\;\mid\;f:\{0,1\}^{n}\rightarrow\{0,1\}\}\subset\mathrm{End}(V^{\otimes(n+1)}$. It is easy to see that $\mathrm{card}(W)=2^{2^{n}}$, and $W$ is $\mathcal{P}$-constructible where $\mathcal{P}$ is a set of $2^{2(n+1)}$ polynomials of degrees bounded by $1$. Now apply Theorem 14. ∎
Proof of Corollary 1.
For $\mathbf{i}=(i_{1},\ldots,i_{q}),1\leq i_{1}<\cdots<i_{q}\leq n+1$, let $A_{\mathbf{i}}\subset\mathrm{End}(V^{\otimes(n+1)})$ be the sub-algebra obtained by tensoring $\mathrm{End}(V)$ in the $i_{j}$-th slots $11\leq i_{1}<\cdots<i_{q}\leq n+$ 1 with $1_{V}$ in the remaining. Thus,
$A_{(1,2,\ldots,q)}=\underbrace{\mathrm{End}(V)\otimes\cdots\otimes\mathrm{End}(V)}_{q}\otimes\overbrace{\mathrm{Id}_{V}\otimes\cdots\otimes\mathrm{Id}_{V}}^{n+1-q}.$
Let $\Delta=\bigcup_{\mathbf{i}=(i_{1},\ldots,i_{q}),1\leq i_{1}<\cdots<i_{q}\leq n+1}A_{\mathbf{i}}$.
Then, $\Delta$ is a union of $\binom{n+1}{q}$ subvarieties of dimension $p=2^{q}\times 2^{q}$, and $D=\deg(\Delta)=\binom{n+1}{q}$.
Moreover, the subset of $A$ consisting of unitary transformations of the form $U_{g,\mathbf{i}}$, where $g$ is a $q$-ary quantum gate and $\mathbf{i}=(i_{1},\ldots,i_{q})$ is a $k$-tuple of distinct indices with $1\leq i_{j}\leq 2^{n+1}$ is contained in $\Delta$ defined above, and using Remark 18, it suffices to bound from below
$\mathrm{rank}_{\mathrm{End}(V^{\otimes(n+1)}),\Delta}(W),$
where $W=\{U_{f}\;\mid\;f:\{0,1\}^{n}\rightarrow\{0,1\}\}\subset\mathrm{End}(V^{\otimes(n+1)})$. The corollary now follows immediately from Theorem 16. ∎
Proof of Theorem 17.
Let
$W^{+}=\bigcup_{f:\{0,1\}^{n}\rightarrow\{0,1\}}U_{f}^{+}.$
Note that $b_{0}(W^{+})=2^{2^{n}}$, since $U_{f}^{+}\cap U_{g}^{+}=\emptyset$ iff $f\neq g$.
It is easy to set that $W^{+}$ is a $\mathcal{P}$-semi-algebraic subset of $A$, for some $\mathcal{P}$ with $\mathrm{card}(\mathcal{P}\leq O(2^{2(n+t)})$ and $\deg(P)\leq 1$.
The rest of the proof is the same as in the proof of Theorem 1, except we use Theorem 15 instead of Theorem 14 to obtain the lower bound in the last step. ∎
Proof of Corollary 2.
The subset $\Delta\subset A$ defined in the proof of Corollary 1 is a union of real algebraic subsets of $A_{\mathbb{R}}$ which are non-singular complete intersections with $\dim\Delta=4^{q}\times 4^{q}$, $\deg\Delta=\binom{n+t+1}{q}$.
Now apply Theorem 17. ∎
5. Open problems and future work
We end with some open problems and suggestions for future work.
  1. 1.
    Is it possible to eliminate the restriction of local constancy of the real dimension in the hypothesis of Theorem 2 ?
  2. 2.
    Is it possible to extend Theorem 1 to a bound on the sum of the Betti numbers of the realizations of all realizable zero-nonzero patterns. The right hand side will definitely have to be larger. For example, the following result is proved in [8].
    Theorem K.
    Let ${\mathcal{P}}\subset\mathrm{C}[X_{1},\ldots,X_{N}]$ be a family of polynomials with #$\#{\mathcal{P}}=s$ and $\deg(P)\leq d$ for all $P\in{\mathcal{P}}$. Then
    $\sum_{\pi\in\{0,1\}^{\mathcal{P}}}b^{BM}({\mathcal{R}}(\pi,\mathrm{C}^{N}))$$\leq$ $\sum_{0\leq\ell\leq N}\left(\binom{s}{N-\ell}\binom{s}{\ell}d^{N}+\sum_{1\leq i\leq\ell}\binom{s}{N-\ell}\binom{s}{\ell-i}d^{O(N^{2})}\right),$
    where $b^{BM}(\cdot)$ denotes the sum of the Borel-Moore Betti numbers.
    Is it possible to prove a similar bound for all realizable zero-nonzero patterns restricted to an algebraic set but where the bound is independent of the ambient dimension $N$ and depends only on the degree of the variety?
  3. 3.
    We believe that the ambient dimension free upper bounds proved in this paper should have other applications (especially in computational complexity theory) than those explored in this paper. We would like to explore such applications.